### v1

# Lifelong Sequential Knowledge Editing without Model Degradation

Abstract
Prior work in parameter-modifying knowledge editing has shown that large-scale sequential editing leads to significant model degradation. In this paper, we study the reasons behind this and scale sequential knowledge editing to 10,000 sequential edits, while maintaining the downstream performance of the original model. We first show that locate-then-edit knowledge editing methods lead to overfitting on the edited facts. We also show that continuous knowledge editing using these methods leads to disproportionate growth in the norm of the edited matrix. We then provide a crucial insight into the inner workings of locate-then-edit methods. We show that norm-growth is a hidden trick employed by these methods that gives larger importance to the output activations produced from the edited layers. With this “importance hacking”, the edited layers provide a much larger contributions to the model’s output. To mitigate these issues, we present ENCORE - Early stopping and Norm-Constrained Robust knowledge Editing. ENCORE controls for overfitting and the disproportionate norm-growth to enable long-term sequential editing, where we are able to perform up to 10,000 sequential edits without loss of downstream performance. ENCORE is also 61% faster than MEMIT and 64% faster than AlphaEdit on Llama3-8B.
Machine Learning, ICML
1 Introduction
(a) MEMIT
(b) AlphaEdit
Figure 1: The continuous growth of norm of edited MLP matrices in LLama3-8B during sequential knowledge editing, as a function of number edits.
Knowledge editing is the task of editing specific facts that a model learns during pretraining in a data and computing efficient manner (De Cao et al., 2021; Yao et al., 2023; Hartvigsen et al., 2024). In this paper, we focus on a category of parameter-modifying knowledge editing methods called “locate-then-edit” methods, which modify only a small subset of the model parameters to add or update knowledge. These methods have been the focus of many recent works (Gu et al., 2024; Gupta et al., 2024a, c; Ma et al., 2024; Kolbeinsson et al., 2024; Fang et al., 2024). While prior work showed knowledge editing methods like ROME (Meng et al., 2022a) and MEMIT (Meng et al., 2022b) lead to a catastophic loss of downstream performance within a few hundred sequential edits (Gu et al., 2024; Gupta et al., 2024b), a recently introduced method called AlphaEdit (Fang et al., 2024) has been able to push this to 3000 edits. In this paper, we aim to enable large-scale sequential knowledge editing without causing model degradation. Towards this goal, we push sequential knowledge editing to what we propose as the next frontier - performing 10,000 sequential knowledge edits without loss of downstream performance.
Our work is based on two main observations. Firstly, we show that existing locate-then-edit knowledge editing methods are prone to overfitting on the edited fact, where the output probability of an edited fact is unusually higher when compared to the confidence with which an unedited model predicts facts. Secondly, we show that sequential knowledge updates made to a model consistently lead to an increase in the norm of the weight matrix being edited, as shown in Figure 1. We show that this norm-growth is a secret trick used by locate-then-edit methods, which we call “importance hacking”. The increasing norm of the edited matrix also increases the norm of the activation vectors produced by the edited layers on average. This allows the outputs produced from the edited layers to have a larger influence on the final output of the model. With this, the edited layers are able to override the information coming from other parts of the model, leading to successful knowledge edits but inadvertently causing a loss of general ability, which might require information coming from other parts of the model.
Based on these observations, we present targeted interventions to overcome the limitations of overfitting on edited facts and disproportionate norm growth. We first formalize locate-then-edit methods as a two-step finetuning process, where the first step uses gradient descent for optimization and the second step directly optimizes the loss using a least-square objective. This way of understanding knowledge editing enable us to apply appropriate interventions depending on the optimization scenario. To mitigate overfitting during the gradient descent process, we propose Most-Probable Early Stopping (MPES) - where we halt gradient descent when edited facts become most probable across all the different contexts used to calculate the loss. MPES reduces overfitting on edited facts and consequently improves editing performance while reducing loss of downstream performance. To counteract the growth of the norm of the edited matrix during sequential knowledge editing, we add a frobenius-norm constraint within the knowledge editing objective of locate-then-edit methods (Meng et al., 2022b). This constraint still allows for a closed-form solution and gives us the benefits of computationally efficient knowledge edits while allowing us to control the increasing norm of edited weight matrices, thus enabling longer and more robust sequential knowledge editing.
Towards large-scale sequential knowledge editing, we present ENCORE - Early stopping and Norm-Constrained Robust knowledge Editing. ENCORE combines MPES with Frobenius-norm constrained objective and enables performing 10,000 sequential knowledge edits while maintaining the downstream performance of the original unedited model. ENCORE significantly outperforms prior locate-then-edit methods like ROME, MEMIT, and AlphaEdit, as verified on GPT2-XL, Llama-2-7B, and Llama-3-8B, and is significantly faster than its predecessors. Specifically, ENCORE is 61% faster than MEMIT and 64% faster than AlphaEdit when editing Llama3-8B.
2 Background and Related Work
In this section, we provide a brief introduction to locate-then-edit knowledge editing methods and present them as a two-step fine-tuning process. For a more detailed introduction to these methods, we refer the reader to prior works (Meng et al., 2022a, b; Gupta et al., 2024c).
Locate-then-edit family of methods like ROME (Meng et al., 2022a) , MEMIT (Meng et al., 2022b) and AlphaEdit (Fang et al., 2024) are used to update facts that can be represented in the form of triplets of the form (subject, relation, object) or $(s,r,o)$. Instead of updating all the weights of a model to incorporate new knowledge, these methods only update certain weight matrices that are most responsible for factual recall (Meng et al., 2022a). The location of an edit within a model is described by a two-dimensional address - (i) an intermediate layer to be edited and (ii) a token from the list of input tokens used to create the target representation. The exact layer to be edited is found using causal tracing (Meng et al., 2022a) or an empirical sweep over all decoder layers of the model (Hase et al., 2024; Yoon et al., 2024). Additionally, updating the second MLP layer in the FFN module of decoder layers has shown optimal knowledge editing performance (Geva et al., 2020; Meng et al., 2022a). This provides the first part of the editing address. Meng et al. (2022a) also showed that using the output representation of the subject token of a sentence produces the best editing results. This provides the second part of the editing address, where the edit is made using the position index of the last token of the subject. We explain this with an example below.
(a) Gradient descent step which finds the target activations for the MLP matrix.
(b) Target activations are used to update the second MLP matrix (in red).
Figure 2: Presenting locate-then-edit knowledge editing methods as a two-step fine-tuning process.
Given a fact to be edited, for example - “The capital of Malaysia is Singapore”, the query phrase for the editing process is “The capital of Malaysia is” and the target phrase is “Singapore”. The first part of the editing address, the exact layer whose second MLP matrix gets edited, is decided before the editing begins. The second part of the editing address is the token index of the last subject token, which in this case would be the last subword-token in “Malasiya”. The intermediate hidden representation of this last subject token is used to make the edit.
Once the editing address has been decided, instead of updating the chosen MLP weight matrix directly using gradient descent, the process of locate-then-edit knowledge editing proceeds in two steps -
  1. 1.
    In the first step (Figure 2(a)), gradient descent is used to find the appropriate activation vector that acts as a target for the weight matrix to be edited. This target activation is found such that the edited fact is generated by the model. In the example, this activation will cause the model to generate “Singapore” in response to the question. Note that in this step, no weights are updated and just an intermediate activation vector is found. This new activation vector now serves as the target for the MLP weight matrix chosen for editing.
  2. 2.
    The weight update happens in the second step of editing (Figure 2(b)), where the MLP matrix is updated with the target activation vector found in the previous step, using a least squares loss function. This loss function tries to preserve the outputs of the MLP matrix for unrelated contexts while generating the target activation when the input corresponds to the query phrase.
Specifically, let $W_{0}$ be the initial weights of the second MLP matrix, which is being edited to $\hat{W}$. $k_{0}$ is used to indicate an input to the MLP matrix representing activation vectors for information we want to preserve from the original model, and $k_{e}$ is input activation vectors representing facts we want to insert into the model. Let $v_{e}$ be the desired target activation vector for the edited MLP matrix found in step 1 of editing using gradient descent. Then the loss function used to update the MLP weight matrix is formulated using least-squares in the form of a preservation-memorization objective (Gupta et al., 2024c):
$\underset{\hat{W}}{\operatorname{argmin}}\hskip 5.0ptL(\hat{W})\hskip 10.0pt\text{where}$ (1)
$L(\hat{W})=\hskip 4.0pt\underbrace{\lambda\sum^{P}_{i=1}\left\|\hat{W}k^{i}_{0}-W_{0}k^{i}_{0}\right\|^{2}_{2}}_{\text{preservation}}+$ $\underbrace{\sum^{B}_{j=1}\left\|\hat{W}k^{j}_{e}-v^{j}_{e}\right\|^{2}_{2}}_{\text{memorization}}$
Since the above objective is linear in the argument, we do not need to use gradient descent for optimization. Thus, locate-then-edit methods can be seen as a unique type of fine-tuning method. Instead of updating the MLP matrix directly using gradient descent on the desired data, the weight update happens in two steps using two different types of objective functions for each step. The first step uses gradient descent whereas the second step uses a closed-form solution.
3 Methods, Models, Datasets and Evaluation
In this paper, we focus on three “locate-then-edit” methods - ROME (Meng et al., 2022a), MEMIT (Meng et al., 2022b) and AlphaEdit (Fang et al., 2024). As discussed in section 2, each of these algorithms use gradient descent to find intermediate target representations followed by a second optimization step which is linear in the argument. ROME uses an equality-constraint to enforce memorization in place of the least square constraint as shown in equation 1, whereas equation 1 represents the objective function used in MEMIT. AlphaEdit adds a null-space projection term in the MEMIT objective.
Prior work has shown model degradation during sequential editing with ROME and MEMIT (Gu et al., 2024; Gupta et al., 2024b), whereas AlphaEdit is a recent method that is able to perform sequential editing for up to 3,000 facts. In this paper, we adopt the experimental setting of AlphaEdit where they perform sequential edits in batches of 100 facts. This means that 100 facts are edited into the model with each weight update, and multiple such updates are performed sequentially. Since ROME only allows for one edit to be made at a time to the model, we used the batched generalization of ROME, called EMMET (Gupta et al., 2024c), which uses the same equality-constraint objective as ROME but generalizes it to the batched editing case.
We evaluate all algorithms on three representative models - GPT2-XL (Radford et al., 2019), Llama2-7B (Touvron et al., 2023) and Llama3-8B (Yoon et al., 2024). All experiments are performed on the CounterFact (Meng et al., 2022a) and zsRE (Mitchell et al., 2021) datasets, which are standard knowledge editing datasets. We present the results for Llama2-7B and Llama3-8B on CounterFact dataset in the main paper and present the remaining results in the appendix due to space constraints.
In this paper, we evaluate the editing algorithms along two dimensions - editing performance and downstream performance. The editing performance evaluates the success of the knowledge editing algorithm in making successful edits, while downstream performance evaluates the extent of model degradation following prior work (Fang et al., 2024; Gupta et al., 2024b; Gu et al., 2024).
Table 1: Comparison between prediction probabilities of edited facts versus facts that a model knows through pretraining. MPES results in more natural prediction probabilities while being 39% - 76% faster than other methods.
Method Model Original Fact Prob Edited Fact Prob w/o MPES Edited Fact Prob w/ MPES Time Per Edit w/o MPES (s) Time Per Edit w/ MPES (s)
EMMET GPT2-XL 0.39 0.98 0.47 1.26 0.63 ($\downarrow$ 50%)
Llama2-7B 0.52 0.92 0.59 6.61 1.86 ($\downarrow$ 71%)
Llama3-8B 0.49 0.99 0.61 7.65 1.86 ($\downarrow$ 76%)
MEMIT GPT2-XL 0.39 0.65 0.14 1.60 0.97 ($\downarrow$ 39%)
Llama2-7B 0.52 0.75 0.41 4.84 2.79 ($\downarrow$ 42%)
Llama3-8B 0.49 0.77 0.41 8.71 3.31 ($\downarrow$ 61%)
AlphaEdit GPT2-XL 0.39 0.98 0.39 1.49 0.83 ($\downarrow$ 44%)
Llama2-7B 0.52 0.84 0.37 5.89 2.69 ($\downarrow$ 54%)
Llama3-8B 0.49 0.75 0.33 9.44 3.43 ($\downarrow$ 63%)
Knowledge Editing Metrics:
To evaluate editing performance, we use five standard knowledge editing metrics (Meng et al., 2022a). (i) Efficacy Score (ES), which measures if an edit was successfully made, (ii) Paraphrase Score (PS), which measures if the model is able to recall edited facts in different scenarios, (iii) Neighborhood Score (NS), which measures if edited facts disturbs unrelated knowledge, (iv) Overall Score (S), which is the harmonic mean of ES, PS and NS, and (v) Generation Entropy (GE), which measures the fluency of a model. A detailed explanation of these metrics is given in Appendix A.
The editing metrics for each model are calculated after making all 10,000 sequential edits (Fang et al., 2024). This approach ensures that the metrics capture both the success of the edits of the latest batch of facts as well as facts that were edited in the past.
Downstream Performance Metrics:
Following prior work (Fang et al., 2024; Gupta et al., 2024b), we measure downstream performance during knowledge editing using 6 tasks - massive multitask language understanding (MMLU) (Hendrycks et al., 2020), natural language inference (NLI, RTE) (Dagan et al., 2005; Haim et al., 2006; Giampiccolo et al., 2007), sentiment analysis (SST2) (Socher et al., 2013), paraphrase detection (MRPC) (Dolan & Brockett, 2005), and linguistic acceptability classification (CoLA) (Warstadt et al., 2019). We measure the downstream performance of models every 1000 edits, following the experimental settings of Fang et al. (2024). More details about experimental details can be found in Appendix J.
4 Overfitting during Knowledge Editing
As discussed in section 2, locate-then-edit methods can be seen as a two-step fine tuning process. The first step uses gradient descent to create intermediate target representations. The gradient descent step minimizes average cross-entropy loss for predicting the target fact over the query phrase augmented by ‘$N$’ random prefixes. The random prefixes are supposed to represent different contexts in which the edited fact can be recalled, thus aiming to create a more general query representation. The average cross-entropy loss over all contexts is optimized as shown:
$L(\theta)=\frac{1}{N}\sum^{j=N}_{j=1}-log\mathcal{P}_{\theta}[o^{*}|x_{j}+p]$ (2)
Here $p$ represents the query phrase input to the model, $o^{*}$ represents the target fact to be edited into the model and $x_{j}$ represent random prefixes added to the query phrase.
(a) MEMIT
(b) AlphaEdit
(c) EMMET
Figure 3: Average downstream performance measured over 6 tasks (sec 2) for MEMIT, AlphaEdit and EMMET. We see that MPES is able to delay loss of downstream performance for both Llama3-8B and Llama2-7B with additional gains in efficiencey.
Table 2: Sequential knowledge editing performance after 10,000 edits for different algorithms. We see that using MPES improves editing metrics across all algorithms and models.
Method Model Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy
MEMIT Llama2-7B 81.04 64.67 60.95 67.859 442.59
Llama3-8B 49.68 49.29 51.31 50.078 373.48
MEMIT Llama2-7B 88.43 70.83 65.86 73.873 542.1
+ MPES Llama3-8B 65.78 57.58 50.25 57.176 560.78
AlphaEdit Llama2-7B 61.1 55.86 53.75 56.74 540.92
Llama3-8B 72.67 63.44 52.9 61.948 465.81
AlphaEdit Llama2-7B 84.15 74.94 62.87 72.933 583.4
+ MPES Llama3-8B 88.43 82.08 56.5 72.832 565.36
EMMET Llama2-7B 94.98 84.05 55.76 74.331 569.24
Llama3-8B 85.03 75.06 49.08 65.995 567.34
EMMET Llama2-7B 96.8 88.4 58.41 77.393 584.28
+ MPES Llama3-8B 92.96 86.12 55.74 74.424 564.82
We observe that when the gradient descent process stops, the intermediate target representations are over-optimized on the edited facts and make the probability of the target fact unusually high. This can be seen in Table 1, where the edited facts are predicted with a probability much closer to 1 compared to the probability with which an unedited model recalls a fact. This phenomenon was also observed by Zhang et al. (2024). More details on the experimental settings can be found in Appendix B.
As seen in Table 1, the probability with which the unedited GPT2-XL, Llama2-7B, and Llama3-8B models recall a fact are 39%, 52% and 49% respectively. However, facts that get edited into the model with algorithms like Alphaedit or EMMET are predicted with average probabilities of 98-99%. The vocabulary size of these models is in the tens of thousands and placing such a high probability mass over every edited fact is a result of clear overfitting. We hypothesize that this overfitting of edited facts is one of the reasons why edited models lose their ability to perform downstream tasks when continuously edited using knowledge editing methods.
4.1 MPES and Knowledge Editing
To overcome this overfitting over a small subset of edited facts, we propose a variant of early stopping called “most-probable early stopping” (MPES). Conventionally, early stopping is used during training while monitoring validation loss, where training is halted when the validation loss stops improving. In MPES, we stop the gradient descent process in knowledge editing when the target fact becomes the most probable token for all ‘$q$’ query phrases used for optimization. While the gradient descent loss continues to go down and the target probability continues to increase, the rank of the predicted word does not change further. With MPES, the rank of the predicted word performs the same function as validation loss in conventional early stopping.
Apart from preventing the model from becoming overly optimized towards a specific target fact, using MPES for halting gradient descent also has two other advantages. Firstly, it simplifies monitoring of the gradient descent process and provides a principled approach to stopping gradient descent which is directly tied to the knowledge editing objective, that is, accurately recalling edited facts in a variety of scenarios without overfitting. Secondly, by optimally stopping the gradient descent process, we also improve the efficiency of locate-then-edit algorithms. MPES leads to a 39-76% reduction in editing time per edit depending on the editing algorithm and model used (Table 1).
The average downstream performance of AlphaEdit, MEMIT, and EMMET for Llama2-7B and Llama3-8B over 10,000 sequential edits is shown in Figure 3111For task-wise downstream performance, please refer to Appendix H. Comparing baseline methods with MPES, we see that MPES shows significant improvements in downstream performance for both AlphaEdit and MEMIT. For EMMET, MPES performs on par with the baseline but is much faster, as shown in Table 1. This shows that by halting gradient descent at a critical point where the edited fact is recognized but not overly dominant, MPES helps maintain the model’s general abilities for much longer during sequential editing.
The editing performance using MPES can be seen in Table 2. We see that MPES not only maintains the model’s general abilities for longer but also improves editing performance for all algorithms. This is because as the model over-optimizes for edited facts, it not only loses its general abilities but also the ability to recall facts edited in previous iterations, resulting in lower editing scores. This also verifies our hypothesis that overfitting over edited facts is one of the reasons behind the loss of model’s general ability.
To summarize, with MPES we present a principled way of stopping the gradient descent step during knowledge editing which results in improved editing performance, delays the loss of downstream performance, and makes current knowledge editing methods much faster.
(a) Alphaedit (Llama3-8B)
(b) MEMIT (Llama3-8B)
Figure 4: Comparison between norm of edited MLP matrices and norm of unedited matrices after 5,000 and 10,000 sequential edits.
5 Norm Growth during Sequential Knowledge Editing
Past work has shown that sequential knowledge editing can lead to an increase in the norm of the edited matrix (Gupta et al., 2024b). This disproportionate growth of norm can be seen in Figure 1 and 4 when using AlphaEdit and MEMIT on Llama3-8B. Both AlphaEdit and MEMIT modify the second MLP matrix of layers 4 to 8. To put the norm growth in perspective with the rest of the model, we plot the norms of the edited matrices after 5k and 10k edits along with the norm of other MLP matrices of the model. We clearly see the disproportionate growth in the norm for the edited layers in Figure 4, where the norm grows to more than 10 times its original value when using MEMIT, and twice the original norm when using AlphaEdit, while the norm of the matrices before and after the edited layers is not changed. The growth of the norm of the edited matrix is a continuous process during knowledge editing. As shown in Figure 1, for not even one edit does the norm of the edited matrix remain constant or decrease during editing. The norm always increases!
But is this growing norm harmful to the model? We answer this question by analyzing the residual stream and reveal a very important property about the inner workings of knowledge editing methods.
5.1 The Secret Mechanics of Knowledge Editing
Due to residual connections in an LLM, the final output of a model can be written as a summation of the outputs of the individual sub-modules. Let $h^{l}$ represent the intermediate hidden state vectors between each decoder layer. Then, the computations within a layer of a general decoder-only LLMs proceed as follows:
[…]
$a^{l}$ $=\texttt{Att}(f^{l})$ (4)
[…]
$m^{l}$ $=W^{l}_{proj}\sigma(W^{l}_{fc}g^{l}+b^{l}_{fc})+b_{proj}$ (6)
$h^{l}$ $=h^{l-1}+a^{l}+m^{l}$ (7)
The intermediate hidden vector between each layer, $h^{l}$, is also sometimes referred to as the residual stream. LN1 represents the first LayerNorm (or equivalently RMSNorm for Llama models) that acts just before the attention module and LN2 represents the second LayerNorm just before the MLP module. Att represents the self-attention module in an LLM whereas the action of a traditional MLP module is written in terms of the individual weight matrices ($W_{fc},W_{proj}$). As the vectors computed in the attention and MLP modules get added back to the residual stream at each layer, the residual stream represents a summation of an increasing number of vectors as we go deeper into the model. A non-recursive formula for the output of the transformer just before unembedding is shown below:
$h^{L}=h^{0}+\sum^{i=L}_{i=1}a^{i}+\sum^{i=L}_{i=1}m^{i}$ (8)
Here, $L$ represents the total number of layers in a model and $h^{L}$ represents the residual stream after the final layer. Note that $h^{L}$ is the activation just before the application of the final layernorm and unembedding matrices. Thus, the output vector at the final layer is a summation of the outputs of individual self-attention and MLP sub-modules.
Now, if the norm of the $W_{proj}$ matrix grows as disproportionately as shown in Figures 1 and 4, the norm of the vectors that are produced from those edited MLP sub-modules will also grow. As the norm of a few vectors in the summation in equation 8 grows, these vectors will begin to dominate the sum. Proof for this is shown in Appendix C.1, where we show that if the norm of a vector in a summation grows, the overall sum effectively tends towards that vector.
(a) Average Norm Proportion For Unedited Llama3-8B
(b) Average Norm Proportion for Llama3-8B using Alphaedit
(c) Average Norm Proportion for Llama3-8B using MEMIT
(d) Average Norm Proportion for Llama3-8B using ENCORE
Figure 5: The figure shows the proportion of contribution of activations vectors produced from each sub-module to the residual stream. The edited layers are shown in red. We see that the influence of the output of the edited layers grows dramatically after contiuous editing.
We also show this effect empirically. The growing norm of activation vectors produced by edited layers before and after editing for Llama3-8B can be seen in Figure 5. Figure 5(a) shows the percentage norm for activation vectors produced from each sub-module of the unedited Llama3-8B model, in comparison to the sum of their norms [^2] This is averaged over 30,000 generated tokens (Appendix C).. After editing using AlphaEdit, which edits layers 4-8 in the Llama model for 10,000 sequential edits (edited layers are shown in red color on x-axis), we can see a drastic increase in the percentage norms for vectors produced by the edited layers (Figure 5(b)). For example, the norm of vectors produced by layer 8, which accounted for only 1% of the total norm, after editing using AlphaEdit accounts for 7% of the norm. This effect is even more drastic for MEMIT as seen in Figure 5(c), where the activation vectors produced by layer 8 account for almost 40% of the total norm, and the vectors produced by all edited layers account for 85% of the total.
This gives us a crucial insight into the inner workings of locate-then-edit methods. Apart from optimizing for the preservation-memorization objective in equation 1, these methods also implicitly increase the norms of the edited matrices, and as a consequence, the norm of the activation vectors produced from those matrices also increases. This makes the representation after the final layer ($h^{L}$) be dominated by the output of edited layers, thus giving the edited layers a larger authority over the final representations. This allows edited layers to override the information produced from other parts of the model, consequently leading to the success of these knowledge editing methods.
This finding is central to explaining the inner workings of locate-then-edit methods, and it also explains their successes and failures. This kind of “importance hacking” makes the output of the edited layers more dominant in the final output and is a shortcut taken by these methods to achieve edit success. However, as the norm of the edited matrix continues to grow when we edit it sequentially, suddenly the final representations begin to largely be composed of the outputs of only the edited layers. This makes it impossible for the model to account for the information processed from other sub-modules which might be important to perform other downstream or unrelated tasks. We hypothesize that this is another cause of the loss of downstream performance.
Table 3: Editing performance of ENCORE when compared to baseline MEMIT and modifications using MPES and Norm Constraint.
Method Model Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy
MEMIT Llama2-7B 81.04 64.67 60.95 67.859 442.59
Llama3-8B 49.68 49.29 51.31 50.078 373.48
MEMIT Llama2-7B 88.43 70.83 65.86 73.873 542.1
+ MPES Llama3-8B 65.78 57.58 50.25 57.17 560.78
MEMIT Llama2-7B 90.94 81.31 59.73 74.931 539.58
+ Norm-Const Llama3-8B 85.72 77.08 58.45 71.86 367.46
ENCORE Llama2-7B 92.57 82.64 60.43 76.043 560.16
Llama3-8B 88.77 78.19 60.07 73.707 523.61
5.2 Introducing ENCORE
In the above discussion, we show how growing norm of edited matrices is detrimental to the functioning of edited models and hypothesize that it is the cause of loss of downstream performance. To test our hypothesis, we propose to add an additional term to the preservation-memorization objective to control this norm growth. Thus, we augment the MEMIT objective with a norm-constraint:
$L(\hat{W})=$ $\underbrace{\lambda_{p}\sum^{P}_{i=1}\left\|\hat{W}k^{i}_{0}-W_{0}k^{i}_{0}\right\|^{2}_{2}}_{\text{preservation}}+\underbrace{\sum^{B}_{j=1}\left\|\hat{W}k^{j}_{e}-v^{j}_{e}\right\|^{2}_{2}}_{\text{memorization}}$ (9)
$\hskip 50.0pt+\underbrace{\lambda_{n}\left\|\hat{W}-W_{0}\right\|^{2}_{F}}_{\text{norm-constraint}}$
The original weight matrix is represented by $W_{0}$ and $\hat{W}$ represents the edited matrix. The above objective has a closed-form solution as shown below (proof, Appendix D):
$\hat{W}$ $=W_{0}+\Delta\hskip 10.0pt\text{where}$ (10)
$\Delta=(V_{1}-W_{0}K_{1})$ $K_{1}^{T}(\lambda_{p}K_{0}K_{0}^{T}+K_{1}K_{1}^{T}+\lambda_{n}I)^{-1}$
Editing performance for 10,000 sequential edits using norm-constraint can be seen in Table 3. Since we add the norm-constraint to the MEMIT objective  [^3] While adding norm-constraint works well with the MEMIT objective, using norm constraint with EMMET and AlphaEdit does not lead to such extensive improvements. We discuss this further in Appendix E., we compare the performance of MEMIT and MEMIT+MPES. We see a significant improvement in all editing metrics for both Llama2-7B and Llama3-8B. The downstream performance can be seen in Figure 6, where norm-constraint sustains downstream performance for much longer compared to MEMIT and MPES.
(a) Llama2-7B
(b) Llama3-8B
Figure 6: Average downstream performance for during sequential editing with ENCORE compared to baseline of MEMIT and addition of MPES and Norm-Constraint (NC).
(a) Comparison with unedited layers.
(b) Norm growth as function of edits.
Figure 7: Growth in norm of edited matrices for ENCORE for Llama3-8B.
To achieve our goal of performing 10,000 sequential edits without loss of downstream performance, we combine MPES with the norm-constraint objective in equation 9 and present ENCORE - Early stopping and Norm-Constrained Robust knowledge Editing. With ENCORE, we bring back the classical knowledge editing objective of MEMIT in a stronger, faster, and more robust way. The editing performance for ENCORE can be seen in Table 3 and Figure 6. For Llama2-7B, we see that maximum improvement in downstream performance (Figure 6(a)) comes from the norm constraint and ENCORE performs equally well. But ENCORE is not only much faster, but also has a better overall editing performance (Table 3). For Llama3-8B (Figure 6(b)), we see that the combination of MPES and norm-constraint is crucial for sustained downstream performance.
ENCORE also controls the growing norm of edited matrices, as shown in Figure 7. As a consequence, the activation norms for ENCORE have a very similar distribution to the original model, which is shown in Figure 5(d). Thus, ENCORE is able to perform large-scale knowledge editing without taking the shortcut of importance hacking, which also prevents the loss of downstream performance. With ENCORE, we combine the insights presented in this paper about overfitting and implicit importance hacking in locate-then-edit methods and enable long-term sequential knowledge editing. ENCORE is able to consistently sustain downstream evaluation metrics of the original model while going through 10,000 sequential edits. ENCORE is also 61% faster than MEMIT and 64% faster than AlphaEdit for Llama3-8B [^4] ENCORE takes 1.01 seconds per edit for GPT2-XL, 3.33 seconds for Llama2-7B and 3.35 seconds for Llama3-8B..
6 Conclusion
In this paper we present two major drawbacks of existing knowledge editing methods - (i) overfitting on edited facts and (ii) continuous and disproportionate norm-growth of edited matrix. We also present insights into the inner workings of locate-then-edit methods and show that they achieve edit success using the short-cut of norm-increase, which increases the importance of the output of edited matrices. We then present ENCORE, a faster and stronger knowledge editing method compared to its predecessors. ENCORE is able to perform 10,000 sequential knowledge edits without noticeable loss of downstream performance in about 40% of the time taken by other algorithms for Llama3-8B. While many recent works have shown limitations of knowledge editing methods at scale, we show that with a better understanding and appropriate regularizations, these methods can indeed be scaled. We hope this brings further research and excitement into the field of knowledge editing.
Impact Statement
This paper advances the field of knowledge editing in LLMs by enabling scalable, updates to a model’s stored knowledge without significant model degradation. Such techniques could be beneficial for multiple practical and societal uses, including quick updating of time-sensitive facts such as changing a model’s knowledge following natural disasters or newly published research.
However, as with any technology, there are potential risks with this method. The ability to efficiently overwrite and retain specific facts raises concerns about malicious editing or the injection of wrong information. Further research on robust governance and the use of these frameworks is still needed. Overall, we believe this research can lead to more responsive, up-to-date, and ethically aligned LLMs provided that careful attention is paid to responsible use.
Appendix A Knowledge Editing Metrics
A more detailed explanation of the knowledge editing metrics used in this paper is below:
  1. 1.
    Efficacy Score (ES): assesses whether an edit has been successful. It is calculated as the percentage of edits where $P(\text{new fact})>P(\text{old fact})$ when evaluated on paraphrases of the query prompt.
  2. 2.
    Paraphrase Score (PS): measures the model’s ability to generalize after an edit. Specifically, it is the percentage of edits where $P(\text{new fact})>P(\text{old fact})$ for paraphrased versions of the query prompt.
  3. 3.
    Neighborhood Score (NS): evaluates the locality of a model edit by determining whether editing one fact affects other facts stored in the model. It is the percentage of unaffected facts in the neighborhood of the edited fact.
  4. 4.
    Generation Entropy (GE): measures the fluency of the model’s text generation post-edit. GE is computed as the weighted average of bi-gram and tri-gram entropies in the text generated by the edited model. A lower GE indicates repetitive text generation, a common failure mode (Meng et al., 2022a; Gupta et al., 2024b).
  5. 5.
    Score (S): introduced by (Meng et al., 2022a), this composite metric combines edit success, generalization, and locality into a single score. It is calculated as the harmonic mean of the Efficacy Score (ES), Paraphrase Score (PS), and Neighborhood Score (NS).
Appendix B Experimental Detail on Overfitting During Knowledge Editing
For each method and model, we conducted three different experiments:
  1. 1.
    Unedited fact recall probability - In this case, we calculate the average probability with which a fact is recalled by the unedited/original model. These are the facts that the model learnt through its pre-training. The model is asked questions from the CounterFact dataset, and we average the probability with which the model predicts the fact correctly.
  2. 2.
    Edited fact probability without MPES - In this case, we evaluate the probability with which a model recalls a fact that is edited into the model. This is the standard baseline case without MPES.
  3. 3.
    Edited fact probability WITH MPES - Here we also evaluate the average probability by which a model recalls an edited fact. In this case, MPES is used during editing the fact.
In each of the experiments we performed, we passed average numbers over 1000 edited facts with a batch size of 1. We use the CounterFact dataset (Meng et al., 2022a) for all of these experiments. We show the result in table 1. As we can see from the table, the unedited fact token probability is pretty low but once we run the edited fact the probability increases to almost 1 for some cases. MPES brings the probability of fact recall for edited facts down to a more natural value, which prevents the overfitting problem that we present in this paper.
Appendix C Experimental Details on Activation Norm Growth
To assess the impact of the activation vectors generated by the edited layers before and after editing, we conducted an experiment where we edited our model on a total of 10,000 facts using a batch size of 100. Once the model was edited, we evaluated the norm of the activations produced by each layer by passing it through a Wikipedia dataset containing 30,000 examples. For each example, the model performed a one-word prediction task given an input context, and we measured the norm of the activation vectors produced from each layer in the model. We repeated the same process for the unedited model to compare the differences.
For each of the 30,000 examples, we calculated the proportion of the activation norm at each layer. We then plotted the mean and standard deviation of these proportions for both the edited and unedited models in figures 8 - 13. As shown in our results, the proportion of activation norms for the layers that were edited is significantly higher than their neighboring layers. In fact, some of the edited layers show the highest proportions overall. The edited layers are highlighted with red color on the x-axis.
C.1 Theoretical Analaysis of Growth of Vector Norm in a Summation
We want to understand the effect of excessive increase in the norm of a vector in a sum of vectors. First, let’s start with an easy example where we have a summation of two vectors, $\mathbb{s}=\mathbb{a}+\mathbb{b}$ and then there is excessive increase in the norm of the first vector, that is $\mathbb{s}=k\mathbb{a}+\mathbb{b}$ where $k$ is some positive scalar. To evaluate the effect of this increase, we analyses the tendencies of the sum $\mathbb{s}$ as k increases.
We first want to understand the norm of $||\mathbb{s}||$. We have the following :
$||\mathbb{s}||^{2}=||k\mathbb{a}+\mathbb{b}||^{2}=k^{2}||\mathbb{a}||^{2}+2k\mathbb{a}\cdot\mathbb{b}+||\mathbb{b}||^{2}$
From this we can clearly see that as $k$ increases the first term quadratic in $k$ will dominate. This means that as $k\to\infty$, $||\mathbb{s}||^{2}\to k^{2}||\mathbb{a}||^{2}$, or $||\mathbb{s}||\to k||\mathbb{a}||$, which is the norm of the new vector. Thus, as the norm of one of the vectors in the summation increases, the norm of the summation tends to the norm of that vector with increasing norm.
Next, we look at the tendencies of the orientation of the summation as the norm of one vector increases. Let $\theta$ be the angle between $\mathbb{s}$ and $k\mathbb{a}$. Then, the cosine of the angle between the summation and the new vector $k\mathbb{a}$ is as follows (note that angle between $\mathbb{s}$ and $k\mathbb{a}$ is same as the angle between $\mathbb{s}$ and $\mathbb{a}$):
$\cos{\theta}=\frac{\mathbb{s}\cdot\mathbb{a}}{||\mathbb{s}||||\mathbb{a}||}=\frac{(k\mathbb{a}+\mathbb{b})\cdot\mathbb{a}}{||k\mathbb{a}+\mathbb{b}||||\mathbb{a}||}$
In the limit of $k\to\infty$, $||\mathbb{s}||\to k||\mathbb{a}||$ as shown above. Thus, the cosine expression in the limit can be written as:
$\underset{k\to\infty}{\cos{\theta}}=\frac{(k\mathbb{a}+\mathbb{b})\cdot\mathbb{a}}{k||\mathbb{a}||^{2}}=\frac{k\mathbb{a}\cdot\mathbb{a}}{k||\mathbb{a}||^{2}}+\frac{\mathbb{b}\cdot\mathbb{a}}{k||\mathbb{a}||^{2}}$
$=1+\frac{\mathbb{b}\cdot\mathbb{a}}{k||\mathbb{a}||^{2}}$
Thus, as $k\to\infty$, the cosine of angle between the sum and the vector tends to 1, or the angle between the summation and the vector tends to zero. This shows that as the norm of a vector in the summation continues to increase, the both the norm and the orientation of the summation tends towards the vector with increasing norm.
Finally, the same proof can be generalied to a summation of multiple vectors, where $\mathbb{b}$ represents the sum of the remaining vectors.
(a) Average Norm Proportion For Unedited GPT2-XL
(b) Average Norm Proportion for GPT2-XL using Alphaedit
(c) Average Norm Proportion for GPT2-XL using EMMET
(d) Average Norm Proportion for GPT2-XL using MEMIT
Figure 8: Activation norm growth for GPT2-XL.
(a) Average Norm Proportion For Unedited Llama2-7B
(b) Average Norm Proportion for Llama2-7B using Alphaedit
(c) Average Norm Proportion for Llama2-7B using EMMET
(d) Average Norm Proportion for Llama2-7B using MEMIT
Figure 9: Activation norm growth for Llama2-7B.
(a) Average Norm Proportion For Unedited Llama3-8B
(b) Average Norm Proportion for Llama3-8B using Alphaedit
(c) Average Norm Proportion for Llama3-8B using EMMET
(d) Average Norm Proportion for Llama3-8B using MEMIT
Figure 10: Activation norm growth for Llama3-8B.
(a) Average Norm Proportion For GPT2-XL using MEMIT
(b) Average Norm Proportion for GPT2-XL using Norm Constraint
(c) Average Norm Proportion for GPT2-XL using ENCORE
Figure 11: Activation norm growth for GPT2-XL using Norm Constraint and ENCORE.
(a) Average Norm Proportion For Llama2-7B using MEMIT
(b) Average Norm Proportion for Llama2-7B using Norm Constraint
(c) Average Norm Proportion for Llama2-7B using ENCORE
Figure 12: Activation norm growth for Llama2-7B using Norm Constraint and ENCORE.
(a) Average Norm Proportion For Llama3-8B using MEMIT
(b) Average Norm Proportion for Llama3-8B using Norm Constraint
(c) Average Norm Proportion for Llama3-8B using ENCORE
Figure 13: Activation norm growth for Llama3-8B using Norm Constraint and ENCORE.
Appendix D Proof for ENCORE Objective
First we have that we can write the equation 9 in term of matrix form where we can stack the $k_{0}^{i}$. Specifically, we define $K_{0}=[k_{0}^{1}\;|k_{0}^{2}\;|\cdots\;|k_{0}^{P}]$, $K_{1}=[k_{e}^{1}\;|k_{e}^{2}\;|\cdots\;|k_{e}^{B}]$ and $V_{1}=[v_{e}^{1}\;|v_{e}^{2}\;|\cdots\;|v_{e}^{B}]$
column wise and instead the L2 norm will become the frobenius norm and we have that
$\lambda_{p}\left\|\hat{W}K_{0}-W_{0}K_{0}\right\|^{2}_{F}+\left\|\hat{W}K_{1}-V_{1}\right\|^{2}_{F}+\lambda_{n}\left\|\hat{W}-W_{0}\right\|^{2}_{F}$
We can differentiate this expression with respect to $\hat{W}$ and let it equal to 0, we get the following
$\lambda_{p}\hat{W}K_{0}K_{0}^{T}-\lambda_{p}W_{0}K_{0}K_{0}^{T}+\hat{W}K_{1}K_{1}^{T}$
$-V_{1}K_{1}^{T}+\lambda_{n}\hat{W}-\lambda_{n}W_{0}=0$
$\lambda_{p}\hat{W}K_{0}K_{0}^{T}+\hat{W}K_{1}K_{1}^{T}+\lambda_{n}\hat{W}$
$=\lambda_{p}W_{0}K_{0}K_{0}^{T}+V_{1}K_{1}^{T}+\lambda_{n}W_{0}$
Since $\hat{W}=W_{0}+\Delta$ we have the following
$\lambda_{p}(W_{0}+\Delta)K_{0}K_{0}^{T}+(W_{0}+\Delta)K_{1}K_{1}^{T}+\lambda_{n}(W_{0}+\Delta)$
$=\lambda_{p}W_{0}K_{0}K_{0}^{T}+V_{1}K_{1}^{T}+\lambda_{n}W_{0}$
$\lambda_{p}W_{0}K_{0}K_{0}^{T}+\lambda_{p}\Delta K_{0}K_{0}^{T}+W_{0}K_{1}K_{1}^{T}+\Delta K_{1}K_{1}^{T}$
$+\lambda_{n}(W_{0}+\Delta)=\lambda_{p}W_{0}K_{0}K_{0}^{T}+V_{1}K_{1}^{T}+\lambda_{n}W_{0}$
$\Delta(\lambda_{p}K_{0}K_{0}^{T}+K_{1}K_{1}^{T}+\lambda_{n}I)=(V_{1}-W_{0}K_{1})K_{1}^{T}$
$\Delta=(V_{1}-W_{0}K_{1})K_{1}^{T}(\lambda_{p}K_{0}K_{0}^{T}+K_{1}K_{1}^{T}+\lambda_{n}I)^{-1}$
Table 5: Editing performance for EMMET with Llama2-7B with norm constraint.
Method Model Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy
EMMET baseline Llama2-7B 93.85 87.32 58.07 76.281 579.79
EMMET best-editing Llama2-7B 94.24 87.2 54.36 74.12 566.08
EMMET best-downstream Llama2-7B 90.35 85.97 53.65 72.57 541.16
Appendix E Norm-Constraint with Other Objectives
While theoretically norm-constraint can be added to both EMMET and AlphaEdit objectives, it doesn’t work well with these objectives. AlphaEdit already has a norm constraint in its objective which is still not able to alleviate norm growth as shown in the main paper. Moreover, when norm-constraint is used with the EMMET objective, we see no appreciable improvements in downstream performance but incur a penalty in editing scores.
We believe that this is the nature of how norm-constraint works. When the norm-constraint is added, it penalizes change in weights from the original weight values. One way to prevent loss of downstream performance is to set the hyperparameter for norm-constraint very high, in which case we will never lose downstream performance. But in this scenario, we will also not be able to edit the model. In the other case, if we don’t restrict norm growth, the editor becomes free to increase the norm of edited matrices, which causes anomalous behaviors as discussed in the main paper. Thus, the norm constraint always comes at the trade-off of editability.
For the MEMIT objective, the norm-constraint helps both editing and downstream performance and the algorithm is best able to make use of this constraint. However, this is not the case for other algorithms. The theoretical reasons why this happens are beyond the scope of this work. But we provide evidence for this trade-off empirically for EMMET below.
Table 5 shows the editing performance of the EMMET baseline without any norm-constraint compared against two cases where norm-constraint is added to its objective. In one case, we present the best editing score that can be achieved when a non-zero norm-constrain hyperparameter is added to the EMMET objective ($\lambda_{n}=20$). In this case, we clearly see that even the best editing score with norm constraint is worse than the baseline EMMET. Figure 14(b) shows the downstream performance for this case, which is even worse than the baseline (Figure 14(a)).
Secondly, in Table 5 we also show the editing score for the hyperparameter that achieves the best downstream performance with non-zero norm-constrain hyperparameter in the EMMET objective ($\lambda_{n}=100$). We clearly see a significant drop from baseline editing performance across all metrics. The downstream performance of this case is shown in Figure 14(c), which is not much different from the baseline (Figure 14(a)). Thus, we show that the norm-constraint does not work well with the EMMET objective.
(a) EMMET baseline
(b) Downstream performance for best editing score with norm constraint.
(c) Downstream peformance for best the scenario where we get best downstream performance with norm constraint.
Figure 14: Comparison between the effect of adding norm-constrain the EMMET objective. We see no appreciable improvement in downstream performace but incur loss in editing performance.
Appendix F Editing Hyperparameters and Hardware Details
Here we present the hyperparameters that we find and used for the MPES, Norm constraint, and ENCORE. Tables 6, 7 and 8 show the hyperparameters for CounterFact dataset and tables 9, 10 and 11 show the hyperparameters for the zsRE dataset. Additionally, there is one more hyperparameter that requires tuning, as stopping immediately when the target token becomes the most probable may not always be optimal. We define this hyperparameter as the probability cutoff, which determines how many additional steps we take before stopping. Specifically, a cutoff of $+n$ means that we continue for $n$ more steps after the target token first becomes the most probable.
All experiments in this paper are done on NVIDIA A6000, including experiments where editing speeds of different methods are timed.
Table 6: Hyperparameters for different algorithms with MPES on CouterFact dataset
Method Model $\lambda_{p}$ Probability Cut Off
EMMET + MPES GPT2-XL 10,000 +1
Llama2-7B 15,000 +0
Llama3-8B 15,000 +0
AlphaEdit + MPEs GPT2-XL 20,000 +1
Llama2-7B 15,000 +0
Llama3-8B 15,000 +0
MEMIT + MPEs GPT2-XL 20,000 +2
Llama2-7B 15,000 +1
Llama3-8B 15,000 +2
Table 7: Hyperparameters for Norm Constraint on CouterFact dataset
Method Model $\lambda_{p}$ $\lambda_{n}$
Norm Constraint GPT2-XL 20,000 10
Llama2-7B 15,000 10
Llama3-8B 15,000 20
Table 8: Hyperparameters for ENCORE on CouterFact dataset
Method Model $\lambda_{p}$ $\lambda_{n}$ Probability Cut Off
ENCORE GPT2-XL 20,000 10 +3
Llama2-7B 15,000 10 +2
Llama3-8B 15,000 20 +1
Table 9: Hyperparameters for different algorithms with MPES on zsRE dataset
Method Model $\lambda_{p}$ Probability Cut Off
EMMET + MPES GPT2-XL 10,000 +1
Llama2-7B 15,000 +1
Llama3-8B 15,000 +2
AlphaEdit + MPEs GPT2-XL 20,000 +1
Llama2-7B 15,000 +1
Llama3-8B 15,000 +3
MEMIT + MPEs GPT2-XL 20,000 +5
Llama2-7B 15,000 +1
Llama3-8B 15,000 +4
Table 10: Hyperparameters for Norm Constraint on zsRE dataset
Method Model $\lambda_{p}$ $\lambda_{n}$
Norm Constraint GPT2-XL 20,000 40
Llama2-7B 15,000 10
Llama3-8B 15,000 20
Table 11: Hyperparameters for ENCORE on zsRE dataset
Method Model $\lambda_{p}$ $\lambda_{n}$ Probability Cut Off
ENCORE GPT2-XL 20,000 40 +4
Llama2-7B 15,000 10 +2
Llama3-8B 15,000 10 +3
Appendix G Norm growth Result
In this section we present the observation of norm growth during editing for different methods (including Norm Constraint and ENCORE) with different models. Figures 15 and 16(c) show the result for different methods for GPT2-XL. Figures 17 and 18(c) show the result for Llama2-7B. Lastly figures 19 and 20(c) show the result for Llama3-8B.
(a) Norm growth of GPT2-XL using Alphaedit
(b) Norm growth of GPT2-XL using EMMET
(c) Norm growth of GPT2-XL using MEMIT
Figure 15: Norm growth of GPT2-XL across different methods
(a) Norm growth of GPT2-XL using MEMIT
(b) Norm growth of GPT2-XL using Norm Constraint
(c) Norm growth of GPT2-XL using ENCORE
Figure 16: Norm growth of GPT2-XL using Norm Constraint and ENCORE
(a) Norm growth of Llama2-7B using Alphaedit
(b) Norm growth of Llama2-7B using EMMET
(c) Norm growth of Llama2-7B using MEMIT
Figure 17: Norm growth of Llama2-7B across different methods
(a) Norm growth of Llama2-7B using MEMIT
(b) Norm growth of Llama2-7B using Norm Constraint
(c) Norm growth of Llama2-7B using ENCORE
Figure 18: Norm growth of Llama2-7B using Norm Constraint and ENCORE
(a) Norm growth of Llama3-8B using Alphaedit
(b) Norm growth of Llama3-8B using EMMET
(c) Norm growth of Llama3-8B using MEMIT
Figure 19: Norm growth of Llama3-8B across different methods
(a) Norm growth of Llama3-8B using MEMIT
(b) Norm growth of Llama3-8B using Norm Constraint
(c) Norm growth of Llama3-8B using ENCORE
Figure 20: Norm growth of Llama3-8B using Norm Constraint and ENCORE
Appendix H Downstream Performance
In this section, we show the downstream performance for each method and also for different algorithms. Note that the reason why x -axis shows the number of edits to be from 0 to 100 is because it shows a number of gradient updates (or batch index) instead of a number of facts edited. Since editing is done in batches of 100, to achieve 10,000 edits we perform a total of 100 batched updates to the models sequentially. Thus, to get the total number of edits finished until a certain point, the x-axis needs to be multiplied by 100, which is the batch size.
Figures 21(a) - 26(a) show the result for the downstream performance for GPT2-XL on both CounterFact and zsRE datasets. Figures 27 - 32 show the result for Llama2-7B on both datasets. Lastly, figures 33 - 38 show the result for Llama3-8B on both datasets.
(a) Downstream Performance for GPT2-XL using AlphaEdit
(b) Downstream Performance for GPT2-XL using AlphaEdit with MPES
Figure 21: Downstream Performance for GPT2-XL using AlphaEdit and MPES with CounterFact dataset
(a) Downstream Performance for GPT2-XL using EMMET
(b) Downstream Performance for GPT2-XL using EMMET with MPES
Figure 22: Downstream Performance for GPT2-XL using EMMET and MPES with CounterFact dataset
(a) Downstream Performance for GPT2-XL using MEMIT
(b) Downstream Performance for GPT2-XL using MEMIT with MPES
(c) Downstream Performance for GPT2-XL using MEMIT with Norm Constraint
(d) Downstream Performance for GPT2-XL using MEMIT with ENCORE
Figure 23: Downstream Performance for GPT2-XL using MEMIT, MPES, Norm Constraint and ENCORE with CounterFact dataset
(a) Downstream Performance for GPT2-XL using AlphaEdit
(b) Downstream Performance for GPT2-XL using AlphaEdit with MPES
Figure 24: Downstream Performance for GPT2-XL using AlphaEdit and MPES with zsRE dataset
(a) Downstream Performance for GPT2-XL using EMMET
(b) Downstream Performance for GPT2-XL using EMMET with MPES
Figure 25: Downstream Performance for GPT2-XL using EMMET and MPES with zsRE dataset
(a) Downstream Performance for GPT2-XL using MEMIT
(b) Downstream Performance for GPT2-XL using MEMIT with MPES
(c) Downstream Performance for GPT2-XL using MEMIT with Norm Constraint
(d) Downstream Performance for GPT2-XL using MEMIT with ENCORE
Figure 26: Downstream Performance for GPT2-XL using MEMIT, MPES, Norm Constraint and ENCORE with zsRE dataset
(a) Downstream Performance for Llama2-7B using AlphaEdit
(b) Downstream Performance for Llama2-7B using AlphaEdit with MPES
Figure 27: Downstream Performance for Llama2-7B using AlphaEdit and MPES with CounterFact dataset
(a) Downstream Performance for Llama2-7B using EMMET
(b) Downstream Performance for Llama2-7B using EMMET with MPES
Figure 28: Downstream Performance for Llama2-7B using EMMET and MPES with CounterFact dataset
(a) Downstream Performance for Llama2-7B using MEMIT
(b) Downstream Performance for Llama2-7B using MEMIT with MPES
(c) Downstream Performance for Llama2-7B using MEMIT with Norm Constraint
(d) Downstream Performance for Llama2-7B using MEMIT with ENCORE
Figure 29: Downstream Performance for Llama2-7B using MEMIT, MPES, Norm Constraint and ENCORE with CounterFact dataset
(a) Downstream Performance for Llama2-7B using AlphaEdit
(b) Downstream Performance for Llama2-7B using AlphaEdit with MPES
Figure 30: Downstream Performance for Llama2-7B using AlphaEdit and MPES with zsRE dataset
(a) Downstream Performance for Llama2-7B using EMMET
(b) Downstream Performance for Llama2-7B using EMMET with MPES
Figure 31: Downstream Performance for Llama2-7B using EMMET and MPES with zsRE dataset
(a) Downstream Performance for Llama2-7B using MEMIT
(b) Downstream Performance for Llama2-7B using MEMIT with MPES
(c) Downstream Performance for Llama2-7B using MEMIT with Norm Constraint
(d) Downstream Performance for Llama2-7B using MEMIT with ENCORE
Figure 32: Downstream Performance for Llama2-7B using MEMIT, MPES, Norm Constraint and ENCORE with zsRE dataset
(a) Downstream Performance for Llama3-8B using AlphaEdit
(b) Downstream Performance for Llama3-8B using AlphaEdit with MPES
Figure 33: Downstream Performance for Llama3-8B using AlphaEdit and MPES with CounterFact dataset
(a) Downstream Performance for Llama3-8B using EMMET
(b) Downstream Performance for Llama3-8B using EMMET with MPES
Figure 34: Downstream Performance for Llama3-8B using EMMET and MPES with CounterFact dataset
(a) Downstream Performance for Llama3-8B using MEMIT
(b) Downstream Performance for Llama3-8B using MEMIT with MPES
(c) Downstream Performance for Llama3-8B using MEMIT with Norm Constraint
(d) Downstream Performance for Llama3-8B using MEMIT with ENCORE
Figure 35: Downstream Performance for Llama3-8B using MEMIT, MPES, Norm Constraint and ENCORE with CounterFact dataset
(a) Downstream Performance for Llama3-8B using AlphaEdit
(b) Downstream Performance for Llama3-8B using AlphaEdit with MPES
Figure 36: Downstream Performance for Llama3-8B using AlphaEdit and MPES with zsRE dataset
(a) Downstream Performance for Llama3-8B using EMMET
(b) Downstream Performance for Llama3-8B using EMMET with MPES
Figure 37: Downstream Performance for Llama3-8B using EMMET and MPES with zsRE dataset
(a) Downstream Performance for Llama3-8B using MEMIT
(b) Downstream Performance for Llama3-8B using MEMIT with MPES
(c) Downstream Performance for Llama3-8B using MEMIT with Norm Constraint
(d) Downstream Performance for Llama3-8B using MEMIT with ENCORE
Figure 38: Downstream Performance for Llama3-8B using MEMIT, MPES, Norm Constraint and ENCORE with zsRE dataset
Appendix I Editing Performance on zsRE dataset
Tables 12 - 14 show the editing scores for the sequential editing experiments on zsRE.
Table 12: Editing performance for GPT2-XL on zsre dataset
Model Method Edit Score Paraphrase Score Neighborhood Score
GPT2-XL EMMET 39.71 27.17 9.78
EMMET+MPES 53.55 39.42 16.81
AlphaEdit 42.1 33.61 14.61
AlphaEdit + MPES 54.99 43.18 18.4
MEMIT 74.6 61.77 22.4
MEMIT + MPES 75.09 61.58 23.37
Norm-Constraint 74.51 61.9 23.39
ENCORE 74.46 61.79 23.41
Table 13: Editing performance for Llama2-7B on zsre dataset
Model Method Edit Score Paraphrase Score Neighborhood Score
Llama2-7B EMMET 75.42 69.69 33.89
EMMET+MPES 84.07 76.9 41.88
AlphaEdit 83.77 77.12 41.96
AlphaEdit + MPES 83.8 77.64 41.97
MEMIT 79.49 74.29 41.8
MEMIT + MPES 83.01 77.45 44.64
Norm-Constraint 88.73 84.05 47.98
ENCORE 89.1 84.28 48.51
Table 14: Editing performance for Llama3-8B on zsre dataset
Model Method Edit Score Paraphrase Score Neighborhood Score
Llama3-8B EMMET 96.97 90.96 45.26
EMMET+MPES 96.75 91.31 46.44
AlphaEdit 89.27 82.19 45.23
AlphaEdit + MPES 93.54 85.93 47.32
MEMIT 96.45 90.3 48.91
MEMIT + MPES 96.85 90.76 47.34
Norm-Constraint 90.4 84.58 49.09
ENCORE 93.15 86.19 49.81
Appendix J Evaluation of Downstream Performance
In this paper, we assess model degradation by measuring downstream performance at regular intervals of edits. Our evaluation suite is wide-ranging and consists of the following 6 tasks – sentiment analysis (SST2) (Socher et al., 2013), paraphrase detection (MRPC) (Dolan & Brockett, 2005), natural language inference (NLI, RTE) (Dagan et al., 2005; Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009), linguistic acceptability classification (CoLA) (Warstadt et al., 2019), and massive multitask language understanding (MMLU) (Hendrycks et al., 2020).
For each task, we created a subset of 100 examples balanced across all multiple-choice options. The models were evaluated on the tasks above, and the accuracy score was measured every 1000 edits. In order to improve the models’ initial performance and achieve meaningful signals, we provided few-shot examples. The few-shot prompt templates used for each task are shown in Figures 39-44.
Review : an exhilarating futuristic thriller-noir , minority report twists the best of technology around a gripping story , delivering a riveting , pulse intensifying escapist adventure of the first order
Sentiment : positive
Review : try as i may , i ca n’t think of a single good reason to see this movie , even though everyone in my group extemporaneously shouted , ‘ thank you ! ’
Sentiment : negative
Review : the film ’s performances are thrilling .
Sentiment : positive
Review : vera ’s technical prowess ends up selling his film short ; he smoothes over hard truths even as he uncovers them .
Sentiment : negative
Review : [input]
Sentiment :
Figure 39: Few shot prompt template used for SST-2
Question: Which expression is equivalent to 4 x 9?
(A) (4x 4) + (4x5)
(B) (4+4) x (4+5)
(C) (4+4)+(4+5)
(D) (4x 4) x (4x5)
Answer: A
Question: A marketing researcher is conducting a survey in a large selling area by contacting a small group of people that is representative of all people in that area. The small, representative group is known as the
(A) population
(B) sample
(C) stratification
(D) universe
Answer: B
Question: A research participant eats half a bowl of M&M candies, and then stops eating. How would a motivation researcher using drive reduction theory explain this participant’s behavior?
(A) Humans are instinctively driven to eat sugar and fat when presented to them.
(B) The Yerkes-Dodson law explains that people will eat food when presented to them, but usually in moderate amounts in order to avoid being perceived as selfish.
(C) The primary drive of hunger motivated the person to eat, and then stop when she/he regained homeostasis.
(D) The research participant was satisfying the second step on the hierarchy of needs: Food needs.
Answer: C
Question: In a deductively valid argument
(A) If all the premises are true, the conclusion must be true
(B) The conclusion has already been stated in its premises
(C) If all the premises are true, the conclusion may still be false
(D) Both A and B
Answer: D
Question: [input]
Answer:
Figure 40: Few shot prompt template used for MMLU
Are the sentences paraphrases of each other.
Sentence 1: Federal regulators have turned from sour to sweet on a proposed \$2.8 billion merger of ice cream giants Nestle Holdings Inc. and Dreyer ’s Grand Ice Cream Inc .
Sentence 2: Federal regulators have changed their minds on a proposed \$2.8 billion merger of ice cream giants Nestle Holdings and Dreyer ’s Grand Ice Cream .
Answer: Yes
Are the sentences paraphrases of each other.
Sentence 1: In the year-ago quarter , the steelmaker recorded a profit of \$16.2 million , or 15 cents per share , on sales of \$1.14 billion .
Sentence 2: In the second quarter last year , AK Steel reported a profit of \$16.2 million , or 15 cents a share .
Answer: No
Are the sentences paraphrases of each other.
Sentence 1: He added : “I ’ve never heard of more reprehensiblebehaviour by a doctor .
Sentence 2: The Harrisons ’ lawyer Paul LiCalsi said : “ I ’ ve never heard of more reprehensible behaviour by a doctor .
Answer: Yes
Are the sentences paraphrases of each other.
Sentence 1: While dioxin levels in the environment were up last year , they have dropped by 75 percent since the 1970s , said Caswell .
Sentence 2: The Institute said dioxin levels in the environment have fallen by as much as 76 percent since the 1970s .
Answer: No
Are the sentences paraphrases of each other.
Sentence 1: [input 1]
Sentence 2: [input 2]
Answer:
Figure 41: Few shot prompt template used for MRPC
Is this sentence linguistically acceptable?
Sentence: The evidence assembled by the prosecution convinced the jury.
Answer: Yes
Is this sentence linguistically acceptable?
Sentence: I live at the place where Route 150 crosses the Hudson River and my dad lives at it too.
Answer: No
Is this sentence linguistically acceptable?
Sentence: The government’s imposition of a fine.
Answer: Yes
Is this sentence linguistically acceptable?
Sentence: Sam gave the ball out of the basket.
Answer: No
Is this sentence linguistically acceptable?
Sentence: [input]
Answer:
Figure 42: Few shot prompt template used for RTE
The town is also home to the Dalai Lama and to more than 10,000 Tibetans living in exile.
Question: The Dalai Lama has been living in exile since 10,000. True or False?
Answer: True
P. Prayong, who like Kevala belongs to the Theravada sect of Buddhism, chose India over other Buddhist majority nations as it is the birthplace of Gautama Buddha.
Question: P. Prayong is a member of Theravada. True or False?
Answer: False
The medical student accused of murdering an erotic masseuse he met on Craigslist is drowning in more than \$100,000 in student loan debt and is so broke he can’t afford to pay an attorney, according to court papers. Philip Markoff, a 23-year-old suspended Boston University medical school student, owes \$130,000 in student loans and does not get money from his parents, leaving him to lean on a taxpayer-funded attorney for his defense, according to a court document in Boston Municipal Court that labels him indigent. Markoff graduated from the State University of New York-Albany and was a second-year medical student at BU.
Question: The medical student Philip Markoff was engaged. True or False?
Answer: True
Traditionally, the Brahui of the Raisani tribe are in charge of the law and order situation through the Pass area. This tribe is still living in present day Balochistan in Pakistan.
Question: The Raisani tribe resides in Pakistan. True or False?
Answer: False
The latest attacks targeted the U-S embassy and a top prosecutor’s office in the Uzbek capital.
Question: [input]. True or False?
Answer:
Figure 43: Few shot prompt template used for CoLA
Turkey is unlikely to become involved in, or allow U.S. forces to use Turkish territory in a Middle East war that does not threaten her territory directly. entails the U.S. to use Turkish military bases.
True or False?
Answer: False
Brooklyn Borough Hall featured a Who’s Who in New York’s literary community during the second annual Brooklyn Book Festival. According to Brooklyn Borough President Marty Markowitz, the borough’s zip code 11215 boasts more authors than anywhere else in the country. It appeared to be the case on Sunday. More than 100 authors were featured at the day-long event, including The Basketball Diaries writer Jim Carroll, former M*A*S*H star Mike Farrell, author and illustrator Mo Willems, Jack Kerouac’s sometime lover and National Book Critics Circle Award recipient Joyce Johnson and PEN American Center President Francine Prose. entails the The Brooklyn Book Festival is held in Brooklyn Borough every year.
True or False?
Answer: True
NASA’s Saturn exploration spacecraft, Cassini , has discovered an atmosphere about the moon Enceladus . This is the first such discovery by Cassini, other than Titan , of the presence of an atmosphere around a Saturn moon. entails the Titan is the fifteenth of Saturn’s known satellites.
True or False?
Answer: False
Dr. Eric Goosby, a pioneer in the fight against AIDS, is President Obama’s choice to run the American effort to combat the disease globally, the White House announced Monday. The President’s Emergency Plan For AIDS Relief, known as Pepfar, was championed by President George W. Bush. It is expected to spend \$48 billion over the next five years and is credited with markedly reducing the disease’s death rate. Its prevention policy has been controversial because of its emphasis on socially conservative methods. With a new administration and a Democratic majority in the House, organizations seeking prevention choices beyond abstinence and fidelity — including a renewed commitment to distributing condoms — are eager to try to rewrite the guidelines. entails the Pepfar is committed to fighting AIDS.
True or False?
Answer: True
[input]
True or False?
Answer:
Figure 44: Few shot prompt template used for NLI

### v2

@@
-Lifelong Sequential Knowledge Editing without Model Degradation
+Lifelong Knowledge Editing requires Better Regularization
 Abstract
-Prior work in parameter-modifying knowledge editing has shown that large-scale sequential editing leads to significant model degradation.
-In this paper, we study the reasons behind this and scale sequential knowledge editing to 10,000 sequential edits, while maintaining the downstream performance of the original model.
-We first show that locate-then-edit knowledge editing methods lead to overfitting on the edited facts.
-We also show that continuous knowledge editing using these methods leads to disproportionate growth in the norm of the edited matrix.
-We then provide a crucial insight into the inner workings of locate-then-edit methods.
-We show that norm-growth is a hidden trick employed by these methods that gives larger importance to the output activations produced from the edited layers.
-With this “importance hacking”, the edited layers provide a much larger contributions to the model’s output.
-To mitigate these issues, we present ENCORE - Early stopping and Norm-Constrained Robust knowledge Editing.
-ENCORE controls for overfitting and the disproportionate norm-growth to enable long-term sequential editing, where we are able to perform up to 10,000 sequential edits without loss of downstream performance.
-ENCORE is also 61% faster than MEMIT and 64% faster than AlphaEdit on Llama3-8B.
-Machine Learning, ICML
+Knowledge editing is a promising way to improve factuality in large language models, but recent studies have shown significant model degradation during sequential editing.
+In this paper, we formalize the popular locate-then-edit methods as a two-step fine-tuning process, allowing us to precisely identify the root cause of this degradation.
+We show that model degradation occurs due to (1) over-optimization of internal activations and (2) continuous norm-growth of edited matrices.
+To mitigate these issues, we introduce two regularization techniques: (1) Most-Probable Early Stopping (MPES) and (2) explicit Frobenius norm-constraint.
+We demonstrate that applying these simple yet effective regularization techniques at key points in the editing process can substantially mitigate model degradation.
+Combining these regularization methods enables scaling locate-then-edit methods to 10,000 edits while reducing editing time by 42-61%.
+These results show that targeted regularization is essential for lifelong knowledge editing.
+Lifelong Knowledge Editing requires Better Regularization
+Akshat Gupta1*, Phudish Preteepamornkul1,2*, Maochuan Lu1 Ahmed Alaa1, Thomas Hartvigsen3, Gopala Anumanchipalli1 1UC Berkeley, 2SCB DataX, 3University of Virginia akshat.gupta@berkeley.edu
 1 Introduction
-(a) MEMIT
-(b) AlphaEdit
-Figure 1: The continuous growth of norm of edited MLP matrices in LLama3-8B during sequential knowledge editing, as a function of number edits.
-Knowledge editing is the task of editing specific facts that a model learns during pretraining in a data and computing efficient manner (De Cao et al., 2021; Yao et al., 2023; Hartvigsen et al., 2024).
-In this paper, we focus on a category of parameter-modifying knowledge editing methods called “locate-then-edit” methods, which modify only a small subset of the model parameters to add or update knowledge.
-These methods have been the focus of many recent works (Gu et al., 2024; Gupta et al., 2024a, c; Ma et al., 2024; Kolbeinsson et al., 2024; Fang et al., 2024).
-While prior work showed knowledge editing methods like ROME (Meng et al., 2022a) and MEMIT (Meng et al., 2022b) lead to a catastophic loss of downstream performance within a few hundred sequential edits (Gu et al., 2024; Gupta et al., 2024b), a recently introduced method called AlphaEdit (Fang et al., 2024) has been able to push this to 3000 edits.
-In this paper, we aim to enable large-scale sequential knowledge editing without causing model degradation.
-Towards this goal, we push sequential knowledge editing to what we propose as the next frontier - performing 10,000 sequential knowledge edits without loss of downstream performance.
-Our work is based on two main observations.
-Firstly, we show that existing locate-then-edit knowledge editing methods are prone to overfitting on the edited fact, where the output probability of an edited fact is unusually higher when compared to the confidence with which an unedited model predicts facts.
-Secondly, we show that sequential knowledge updates made to a model consistently lead to an increase in the norm of the weight matrix being edited, as shown in Figure 1.
-We show that this norm-growth is a secret trick used by locate-then-edit methods, which we call “importance hacking”.
-The increasing norm of the edited matrix also increases the norm of the activation vectors produced by the edited layers on average.
-This allows the outputs produced from the edited layers to have a larger influence on the final output of the model.
-With this, the edited layers are able to override the information coming from other parts of the model, leading to successful knowledge edits but inadvertently causing a loss of general ability, which might require information coming from other parts of the model.
-Based on these observations, we present targeted interventions to overcome the limitations of overfitting on edited facts and disproportionate norm growth.
-We first formalize locate-then-edit methods as a two-step finetuning process, where the first step uses gradient descent for optimization and the second step directly optimizes the loss using a least-square objective.
-This way of understanding knowledge editing enable us to apply appropriate interventions depending on the optimization scenario.
-To mitigate overfitting during the gradient descent process, we propose Most-Probable Early Stopping (MPES) - where we halt gradient descent when edited facts become most probable across all the different contexts used to calculate the loss.
-MPES reduces overfitting on edited facts and consequently improves editing performance while reducing loss of downstream performance.
-To counteract the growth of the norm of the edited matrix during sequential knowledge editing, we add a frobenius-norm constraint within the knowledge editing objective of locate-then-edit methods (Meng et al., 2022b).
-This constraint still allows for a closed-form solution and gives us the benefits of computationally efficient knowledge edits while allowing us to control the increasing norm of edited weight matrices, thus enabling longer and more robust sequential knowledge editing.
-Towards large-scale sequential knowledge editing, we present ENCORE - Early stopping and Norm-Constrained Robust knowledge Editing.
-ENCORE combines MPES with Frobenius-norm constrained objective and enables performing 10,000 sequential knowledge edits while maintaining the downstream performance of the original unedited model.
-ENCORE significantly outperforms prior locate-then-edit methods like ROME, MEMIT, and AlphaEdit, as verified on GPT2-XL, Llama-2-7B, and Llama-3-8B, and is significantly faster than its predecessors.
-Specifically, ENCORE is 61% faster than MEMIT and 64% faster than AlphaEdit when editing Llama3-8B.
-2 Background and Related Work
-In this section, we provide a brief introduction to locate-then-edit knowledge editing methods and present them as a two-step fine-tuning process.
-For a more detailed introduction to these methods, we refer the reader to prior works (Meng et al., 2022a, b; Gupta et al., 2024c).
-Locate-then-edit family of methods like ROME (Meng et al., 2022a) , MEMIT (Meng et al., 2022b) and AlphaEdit (Fang et al., 2024) are used to update facts that can be represented in the form of triplets of the form (subject, relation, object) or $(s,r,o)$.
-Instead of updating all the weights of a model to incorporate new knowledge, these methods only update certain weight matrices that are most responsible for factual recall (Meng et al., 2022a).
-The location of an edit within a model is described by a two-dimensional address - (i) an intermediate layer to be edited and (ii) a token from the list of input tokens used to create the target representation.
-The exact layer to be edited is found using causal tracing (Meng et al., 2022a) or an empirical sweep over all decoder layers of the model (Hase et al., 2024; Yoon et al., 2024).
-Additionally, updating the second MLP layer in the FFN module of decoder layers has shown optimal knowledge editing performance (Geva et al., 2020; Meng et al., 2022a).
-This provides the first part of the editing address.
-Meng et al.
-(2022a) also showed that using the output representation of the subject token of a sentence produces the best editing results.
-This provides the second part of the editing address, where the edit is made using the position index of the last token of the subject.
-We explain this with an example below.
-(a) Gradient descent step which finds the target activations for the MLP matrix.
+Knowledge editing entails editing specific facts that a language model has learned during pre-training in a data and computate efficient manner De Cao et al.
+(2021); Yao et al.
+(2023).
+The most popular editors follow a “locate-then-edit” approach, where the methods aim to locate and update small subsets of a model’s parameters that recall target facts Meng et al.
+(2022a, b); Gupta et al.
+(2024a, c); Ma et al.
+(2024); Fang et al.
+(2024).
+Although these methods have shown strong results when performing singular or small-scale edits, they suffer substantial degradation when applied at scale Gu et al.
+(2024a); Gupta et al.
+(2024b); Thede et al.
+(2025), leaving the problem of lifelong knowledge editing largely unsolved.
+* Equal Contribution
+(a) Norm-growth as function of number of edits
+(b) Layer wise norm post-editing
+Figure 1: The continuous growth of norm of edited MLP matrices in MEMIT Llama3-8B during sequential knowledge editing.
+In this work, we systematically investigate the factors that cause model degradations and propose appropriate regularization methods to mitigate them, enabling large-scale sequential knowledge editing without sacrificing downstream performance.
+We begin by first formalizing locate-then-edit methods as a two-step fine-tuning process (section 4), where the first step uses gradient-descent for activation optimization, and the second step performs weight-update via a least-squares update rule.
+This explicit structure allows us to precisely diagnose the causes of loss of downstream performance and allows us to choose appropriate interventions.
+We show that model degradation occurs during continuous sequential editing because of two reasons.
+Firstly, we show that the gradient-descent step in locate-then-edit methods leads to over-optimization of internal activations, which makes the model predict edited facts with unnaturally high confidence (section 5).
+We mitigate this by proposing Most-Probable Early Stopping (MPES) - a novel variant of early-stopping where we halt gradient-descent steps when edited facts become most probable across all the different contexts used to calculate the loss.
+Secondly, we show that sequential editing leads to disproportionate growth of the Frobenius norm of the edited matrix (Figure 1).
+This allows the outputs produced from the edited layers to have an abnormally larger influence on the final output of the model, inadvertently causing loss of general ability, which might require information coming from other parts of the model.
+To address this, we incorporate a Frobenius norm-constraint into the editing objective of locate-then-edit methods (section 6.2).
+While early stopping and norm-constraints are known regularization methods, their targeted stage-specific application to locate-then-edit methods is non-trivial.
+With our work, we demonstrate that there exists a lack of proper regularization in existing knowledge editing methods and show that explicit regularization at the appropriate stages of editing is essential and critical for scalable and stable knowledge editing.
+The code for our work can be found here - https://github.com/scalable-model-editing/knowledge-editing-regularization
+Our proposed regularization methods mitigate the respective issues of over-optimization of intermediate activation vectors and disproportionate norm-growth, consequently preserving downstream performance for a larger number of edits.
+Finally, we show that combining these two regularization methods allows us to perform up to 10,000 sequential edits while maintaining original downstream performance and making editing 41-62% faster.
+2 Related Work
+Knowledge editing methods can broadly be classified into three types (Yao et al., 2023; Zhang et al., 2024c) - hypernetwork based editing methods (De Cao et al., 2021; Mitchell et al., 2021), where an additional model is trained to predict the edited weights of the model, in-context editing methods (Mitchell et al., 2022; Hartvigsen et al., 2024; Zhong et al., 2023), where edited information is added in-context, and locate-then-edit methods Meng et al.
+(2022a, b); Gupta et al.
+(2024c, a); Fang et al.
+(2024), where the knowledge source is traced to certain MLP layers Hase et al.
+(2024); Yoon et al.
+(2024) before modifying them using a two-step fine-tuning method (section 4).
+This work focuses on performing locate-then-edit type of knowledge editing methods.
+Overfitting During Knowledge Editing.
+Some recent works have studied overfitting as a key challenge of knowledge editing.
+For example, Zhang et al.
+(2024a) identifies that locate-then-edit methods overfit on edited facts.
+They propose to mitigate this by introducing an inference-stage regularization approach called “learn to inference” (LTI) that guides edited models to recall new knowledge by adding multi-stage constraints during optimization.
+With our proposed variant of early-stopping called "most-probable early stopping" (MPES), we present a much simpler way of reducing overfitting.
+We also show that LTI does not scale to multiple edits and leads to immediate model degradation when performing sequential editing.
+Meanwhile, MPES leads to significant improvements in downstream performance while countering overfitting (section 5).
+Norm-Growth and Weight Regularization.
+Prior work has shown that sequential knowledge editing leads to the norm-growth of the updated weight matrices Gupta et al.
+(2024b); Yang et al.
+(2024).
+Ma et al.
+(2025) present a similar observation in the form of growth of condition number of the edited matrices as a potential cause of loss of downstream performance.
+However, none of these works explains why increasing norm of edited matrix is bad for the model.
+In our work, we analyze the effect of norm-growth on internal activations and show how it connects with model degradations (section 6.1).
+We verify our conclusions by adding a norm-constraint in the optimization objective of MEMIT.
+We show that our proposed norm-constraint combines seamlessly with MEMIT’s closed-form solution, unlike prior regularization methods like PRUNE Ma et al.
+(2025) and RECT Gu et al.
+(2024b), which require additional regularization steps post-editing.
+We also show that our methods significantly outperform PRUNE and RECT in preserving downstream performance over a large number of edits.
+3 Methods, Models, and Evaluation
+We adopt the experimental setting of AlphaEdit Fang et al.
+(2024), a recent locate-then-edit method that is able to perform sequential editing for up to 3,000 facts.
+Following their setting, we perform sequential edits in batches of 100 facts.
+This means that 100 facts are edited into the model with each weight update, and multiple such updates are performed sequentially.
+We evaluate all algorithms on three representative models - GPT2-XL Radford et al.
+(2019), Llama2-7B Touvron et al.
+(2023) and Llama3-8B Yoon et al.
+(2024).
+All experiments are performed on the CounterFact Meng et al.
+(2022a) and zsRE Mitchell et al.
+(2021) datasets, which are standard knowledge editing datasets.
+We present the results for Llama2-7B and Llama3-8B on the CounterFact dataset in the main paper and present the remaining results in the appendix due to space constraints.
+In this paper, we evaluate the editing algorithms along two dimensions - editing performance and downstream performance.
+The editing performance evaluates the success of the knowledge editing algorithm in making successful edits, while downstream performance evaluates the extent of model degradation following prior work Fang et al.
+(2024); Gupta et al.
+(2024b); Gu et al.
+(2024a).
+(a) gradient-descent step which finds the target activations for the MLP matrix.
 (b) Target activations are used to update the second MLP matrix (in red).
 Figure 2: Presenting locate-then-edit knowledge editing methods as a two-step fine-tuning process.
-Given a fact to be edited, for example - “The capital of Malaysia is Singapore”, the query phrase for the editing process is “The capital of Malaysia is” and the target phrase is “Singapore”.
+Knowledge Editing Metrics:
+To evaluate editing performance, we use five standard knowledge editing metrics Meng et al.
+(2022a).
+(i) Efficacy Score (ES), which measures if an edit was successfully made, (ii) Paraphrase Score (PS), which measures if the model is able to recall edited facts in different scenarios, (iii) Neighborhood Score (NS), which measures if edited facts disturb unrelated knowledge, (iv) Overall Score (S), which is the harmonic mean of ES, PS, and NS, and (v) Generation Entropy (GE), which measures the fluency of a model.
+A detailed explanation of these metrics is given in Appendix A.
+The editing metrics for each model are calculated after making all 10,000 sequential edits Fang et al.
+(2024).
+This approach ensures that the metrics capture both the success of the edits of the latest batch of facts as well as facts that were edited in the past.
+Downstream Performance Metrics:
+Following prior work Fang et al.
+(2024); Gupta et al.
+(2024b), we measure downstream performance during knowledge editing using 6 tasks - massive multitask language understanding (MMLU) Hendrycks et al.
+(2020), natural language inference (NLI, RTE) Dagan et al.
+(2005); Haim et al.
+(2006); Giampiccolo et al.
+(2007), sentiment analysis (SST2) Socher et al.
+(2013), paraphrase detection (MRPC) Dolan and Brockett (2005), and linguistic acceptability classification (CoLA) Warstadt et al.
+(2019).
+Performance is assessed every 1000 edits, following Fang et al.
+(2024).
+Additional details are provided in Appendix N.
+4 Knowledge Editing as a Two-Step Fine-tuning Process
+“Locate-then-edit" family of methods like ROME Meng et al.
+(2022a) , MEMIT Meng et al.
+(2022b) and AlphaEdit Fang et al.
+(2024) update factual knowledge expressed as (subject, relation, object) or (s,r,o).
+Instead of updating all parameters of a model to incorporate new knowledge, these methods only update certain matrices that are most responsible for fact recall Meng et al.
+(2022a).
+The location of an edit within a model is described by a two-dimensional address - (i) an intermediate layer to be edited and (ii) a token from the list of input tokens used to create the target representation.
+The exact layer to be edited is found via causal tracing Meng et al.
+(2022a) or an empirical sweep across the model layers Hase et al.
+(2024); Yoon et al.
+(2024).
+Only the second MLP matrix in the FFN module of the editing layer is updated (Geva et al., 2020; Meng et al., 2022a), forming the first part of the editing address.
+Meng et al.
+(2022a) also showed that using the output representation of the last subject token produces the best editing results, providing the second part of the address.
+The entire editing process is presented in Figure 2.
+We explain this with an example.
+Given a fact to be edited, for example - “The capital of Malaysia is Singapore", the query phrase for the editing process is “The capital of Malaysia is" and the target phrase is “Singapore".
 The first part of the editing address, the exact layer whose second MLP matrix gets edited, is decided before the editing begins.
-The second part of the editing address is the token index of the last subject token, which in this case would be the last subword-token in “Malasiya”.
+The second part of the editing address is the token index of the last subject token, which in this case would be the last subword-token in “Malasiya".
 The intermediate hidden representation of this last subject token is used to make the edit.
-Once the editing address has been decided, instead of updating the chosen MLP weight matrix directly using gradient descent, the process of locate-then-edit knowledge editing proceeds in two steps -
+Once the editing address has been decided, instead of updating the chosen MLP weight matrix directly using gradient-descent, locate-then-edit knowledge editing proceeds in two steps.
   1. 1.
-    In the first step (Figure 2(a)), gradient descent is used to find the appropriate activation vector that acts as a target for the weight matrix to be edited.
-This target activation is found such that the edited fact is generated by the model.
-In the example, this activation will cause the model to generate “Singapore” in response to the question.
+    In the first step (Figure 2(a)), gradient-descent is used to find the appropriate activation vector that acts as a target for the weight matrix to be edited.
+In the example, the found activation will cause the model to generate “Singapore" in response to the question.
+The loss function for the gradient-descent step is shown in equation 2.
 Note that in this step, no weights are updated and just an intermediate activation vector is found.
-This new activation vector now serves as the target for the MLP weight matrix chosen for editing.
   2. 2.
-    The weight update happens in the second step of editing (Figure 2(b)), where the MLP matrix is updated with the target activation vector found in the previous step, using a least squares loss function.
-This loss function tries to preserve the outputs of the MLP matrix for unrelated contexts while generating the target activation when the input corresponds to the query phrase.
-Specifically, let $W_{0}$ be the initial weights of the second MLP matrix, which is being edited to $\hat{W}$.
-$k_{0}$ is used to indicate an input to the MLP matrix representing activation vectors for information we want to preserve from the original model, and $k_{e}$ is input activation vectors representing facts we want to insert into the model.
-Let $v_{e}$ be the desired target activation vector for the edited MLP matrix found in step 1 of editing using gradient descent.
-Then the loss function used to update the MLP weight matrix is formulated using least-squares in the form of a preservation-memorization objective (Gupta et al., 2024c):
+    The weight update occurs in the second editing step (Figure 2(b)), where the MLP matrix is updated using the target activation vector found previously.
+This update uses a least square loss function, which preserves the MLP outputs for unrelated contexts while generating the target activation when the input corresponds to the query phrase.
+The loss function used to update the MLP weight matrix is formulated using least-squares in the form of a preservation-memorization objective Gupta et al. (2024c):
 $\underset{\hat{W}}{\operatorname{argmin}}\hskip 5.0ptL(\hat{W})\hskip 10.0pt\text{where}$ (1)
 $L(\hat{W})=\hskip 4.0pt\underbrace{\lambda\sum^{P}_{i=1}\left\|\hat{W}k^{i}_{0}-W_{0}k^{i}_{0}\right\|^{2}_{2}}_{\text{preservation}}+$ $\underbrace{\sum^{B}_{j=1}\left\|\hat{W}k^{j}_{e}-v^{j}_{e}\right\|^{2}_{2}}_{\text{memorization}}$
-Since the above objective is linear in the argument, we do not need to use gradient descent for optimization.
-Thus, locate-then-edit methods can be seen as a unique type of fine-tuning method.
-Instead of updating the MLP matrix directly using gradient descent on the desired data, the weight update happens in two steps using two different types of objective functions for each step.
-The first step uses gradient descent whereas the second step uses a closed-form solution.
-3 Methods, Models, Datasets and Evaluation
-In this paper, we focus on three “locate-then-edit” methods - ROME (Meng et al., 2022a), MEMIT (Meng et al., 2022b) and AlphaEdit (Fang et al., 2024).
-As discussed in section 2, each of these algorithms use gradient descent to find intermediate target representations followed by a second optimization step which is linear in the argument.
-ROME uses an equality-constraint to enforce memorization in place of the least square constraint as shown in equation 1, whereas equation 1 represents the objective function used in MEMIT.
-AlphaEdit adds a null-space projection term in the MEMIT objective.
-Prior work has shown model degradation during sequential editing with ROME and MEMIT (Gu et al., 2024; Gupta et al., 2024b), whereas AlphaEdit is a recent method that is able to perform sequential editing for up to 3,000 facts.
-In this paper, we adopt the experimental setting of AlphaEdit where they perform sequential edits in batches of 100 facts.
-This means that 100 facts are edited into the model with each weight update, and multiple such updates are performed sequentially.
-Since ROME only allows for one edit to be made at a time to the model, we used the batched generalization of ROME, called EMMET (Gupta et al., 2024c), which uses the same equality-constraint objective as ROME but generalizes it to the batched editing case.
-We evaluate all algorithms on three representative models - GPT2-XL (Radford et al., 2019), Llama2-7B (Touvron et al., 2023) and Llama3-8B (Yoon et al., 2024).
-All experiments are performed on the CounterFact (Meng et al., 2022a) and zsRE (Mitchell et al., 2021) datasets, which are standard knowledge editing datasets.
-We present the results for Llama2-7B and Llama3-8B on CounterFact dataset in the main paper and present the remaining results in the appendix due to space constraints.
-In this paper, we evaluate the editing algorithms along two dimensions - editing performance and downstream performance.
-The editing performance evaluates the success of the knowledge editing algorithm in making successful edits, while downstream performance evaluates the extent of model degradation following prior work (Fang et al., 2024; Gupta et al., 2024b; Gu et al., 2024).
-Table 1: Comparison between prediction probabilities of edited facts versus facts that a model knows through pretraining.
-MPES results in more natural prediction probabilities while being 39% - 76% faster than other methods.
-Method Model Original Fact Prob Edited Fact Prob w/o MPES Edited Fact Prob w/ MPES Time Per Edit w/o MPES (s) Time Per Edit w/ MPES (s)
-EMMET GPT2-XL 0.39 0.98 0.47 1.26 0.63 ($\downarrow$ 50%)
-Llama2-7B 0.52 0.92 0.59 6.61 1.86 ($\downarrow$ 71%)
-Llama3-8B 0.49 0.99 0.61 7.65 1.86 ($\downarrow$ 76%)
-MEMIT GPT2-XL 0.39 0.65 0.14 1.60 0.97 ($\downarrow$ 39%)
-Llama2-7B 0.52 0.75 0.41 4.84 2.79 ($\downarrow$ 42%)
-Llama3-8B 0.49 0.77 0.41 8.71 3.31 ($\downarrow$ 61%)
-AlphaEdit GPT2-XL 0.39 0.98 0.39 1.49 0.83 ($\downarrow$ 44%)
-Llama2-7B 0.52 0.84 0.37 5.89 2.69 ($\downarrow$ 54%)
-Llama3-8B 0.49 0.75 0.33 9.44 3.43 ($\downarrow$ 63%)
-Knowledge Editing Metrics:
-To evaluate editing performance, we use five standard knowledge editing metrics (Meng et al., 2022a).
-(i) Efficacy Score (ES), which measures if an edit was successfully made, (ii) Paraphrase Score (PS), which measures if the model is able to recall edited facts in different scenarios, (iii) Neighborhood Score (NS), which measures if edited facts disturbs unrelated knowledge, (iv) Overall Score (S), which is the harmonic mean of ES, PS and NS, and (v) Generation Entropy (GE), which measures the fluency of a model.
-A detailed explanation of these metrics is given in Appendix A.
-The editing metrics for each model are calculated after making all 10,000 sequential edits (Fang et al., 2024).
-This approach ensures that the metrics capture both the success of the edits of the latest batch of facts as well as facts that were edited in the past.
-Downstream Performance Metrics:
-Following prior work (Fang et al., 2024; Gupta et al., 2024b), we measure downstream performance during knowledge editing using 6 tasks - massive multitask language understanding (MMLU) (Hendrycks et al., 2020), natural language inference (NLI, RTE) (Dagan et al., 2005; Haim et al., 2006; Giampiccolo et al., 2007), sentiment analysis (SST2) (Socher et al., 2013), paraphrase detection (MRPC) (Dolan & Brockett, 2005), and linguistic acceptability classification (CoLA) (Warstadt et al., 2019).
-We measure the downstream performance of models every 1000 edits, following the experimental settings of Fang et al.
+Specifically, $W_{0}$ represents the initial weights of the second MLP matrix, which is being edited to $\hat{W}$.
+$k_{0}$ represent input to the MLP matrix for information we want to preserve from the original model, and $k_{e}$ is input activation vectors representing facts we want to insert into the model.
+$v_{e}$ is the target activation vector found in step 1 of editing using gradient-descent.
+Since the above objective is linear in the argument, we do not need to use gradient-descent for optimization.
+Thus, locate-then-edit methods can be seen as a unique type of 2-step fine-tuning method.
+Instead of updating the MLP matrix directly using gradient-descent on the desired data, the weight update happens in two steps using two different types of objective functions for each step.
+The first step uses gradient-descent whereas the second step uses a closed-form solution.
+Model Prediction Probability Time / edit (s) Original fact MEMIT MEMIT w/ LTI MEMIT w/ MPES MEMIT MEMIT w/ LTI MEMIT w/ MPES Llama-2 7B 0.52 0.78 0.30 0.45 4.84 8.67 2.79 ($\downarrow$ 42 %) Llama-3 8B 0.49 0.79 0.29 0.41 8.71 9.21 3.31 ($\downarrow$ 61 %)
+Table 1: Comparison between prediction probabilities of edited facts along with editing time when using MPES and LTI Zhang et al. (2024b) with MEMIT.
+5 Over-Optimization of Target Activations
+The gradient-descent step for finding target activations as described in section 4 minimizes average cross-entropy loss for predicting the target fact for the query phrase augmented by ‘$N$’ random prefixes.
+The random prefixes are supposed to represent different contexts in which the edited fact can be recalled, thus aiming to create a more general query representation Meng et al.
+(2022a).
+Let $p$ represent the query-phrase input to the model, $o^{*}$ represent the target fact to be edited into the model and $x_{j}$ represent random prefixes added to the query phrase.
+Then, the gradient-descent step minimizes the following loss function:
+$L(\theta)=\frac{1}{N}\sum^{j=N}_{j=1}-log\mathcal{P}_{\theta}[o^{*}|x_{j}+p],$ (2)
+Method Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy Llama2-7B Llama3-8B Llama2-7B Llama3-8B Llama2-7B Llama3-8B Llama2-7B Llama3-8B Llama2-7B Llama3-8B MEMIT 81.04 49.68 64.67 49.29 60.95 51.31 67.86 50.08 442.59 373.48 MEMIT + LTI 55.34 51.85 53.57 51.23 49.21 49.21 52.58 50.70 562.07 372.08 MEMIT + MPES 93.21 74.40 83.43 66.32 62.25 50.78 77.36 62.23 569.84 466.97 RECT 82.42 63.17 66.84 56.92 67.39 52.89 71.54 57.36 549.35 588.39 RECT + LTI 53.87 51.55 52.17 51.08 50.35 50.43 52.09 51.02 498.97 544.41 RECT + MPES 92.12 66.52 78.06 59.11 67.16 59.44 77.81 61.51 596.62 410.82 PRUNE 70.80 49.38 62.11 49.63 51.86 51.09 60.60 50.02 280.83 340.22 PRUNE + LTI 55.19 49.88 52.98 49.63 50.33 50.19 52.76 49.90 548.23 239.17 PRUNE + MPES 91.48 92.66 83.65 82.20 61.91 64.82 76.85 78.16 568.91 512.83 AlphaEdit 61.10 72.67 55.86 63.44 53.75 52.90 56.74 61.95 540.92 465.81 AlphaEdit + LTI 53.82 53.27 53.03 51.70 49.68 49.72 52.11 51.52 524.11 240.07 AlphaEdit + MPES 84.15 88.43 74.94 82.08 62.87 56.60 72.93 72.83 583.40 565.32
+Table 2: Sequential knowledge-editing performance after 10,000 edits.
+Scores are shown for nine MEMIT-based and AlphaEdit algorithms under two models.
+MPES consistently improves editing metrics across all settings.
+We observe that when the gradient-descent step stops in these algorithms, the intermediate target representations become over-optimized, assigning unusually high probabilities to edited facts.
+More specifically, the average target probability at which the gradient-descent step stops for all locate-then-end algorithms lies between 95-100%.
+When this target representation is used to update the the edited matrix using equation 1, the resultant edits are predicted with abnormally high probabilities.
+This can be seen in Table 1, where the edited facts are predicted with a much larger probability compared to the original “unedited” models.
+More details on the experimental settings can be found in Appendix B.
+This phenomenon was also observed by Zhang et al.
+(2024b).
+5.1 MPES and Knowledge Editing
+As seen in Table 1, the probability with which the unedited Llama2-7B, and Llama3-8B naturally recall a fact are 52% and 49% respectively.
+However, facts that get edited into the model with algorithms like MEMIT are predicted with an average probability of 78-79%.
+To overcome this hyper-optimization over a small subset of edited facts, we propose a variant of early stopping called “most-probable early stopping" (MPES).
+Conventionally, early stopping is used during training while monitoring validation loss, where training is halted when the validation loss stops improving.
+In MPES, we stop the gradient-descent step in knowledge editing when the target fact becomes the most probable token for all ‘$N$’ query phrases used for optimization (equation 2).
+This contrasts with the current stopping criteria in these methods, where gradient-descent is halted after either 20 iterations or upon reaching a loss threshold.
+These criteria leave open the possibility of over-optimization (and hence overfitting, as observed empirically in Table 1), or underfitting.
+Apart from preventing the model from becoming overly optimized towards a specific target fact, using MPES for halting gradient-descent also has two other advantages.
+Firstly, it simplifies monitoring of the gradient-descent process and provides a principled approach to stopping gradient-descent which is directly tied to the knowledge editing objective, that is, accurately recalling edited facts in a variety of scenarios without overfitting.
+Secondly, by optimally stopping the gradient-descent process, we also improve the efficiency of locate-then-edit algorithms.
+MPES reduces the average edit time by 42-61% compared to the standard MEMIT.
+(a) Llama2-7B
+(b) Llama3-8B
+Figure 3: Downstream evaluation when comparing MPES (our method) with LTI for regularizing over-optimization of target activations during knowledge editing.
+We compare MPES with LTI Zhang et al.
+(2024b), which prevents over-optimization during the gradient-descent step using multiple additional loss functions that also increase the average time per edit (LTI takes approximately three times as long as MPES as shown in Table 1).
+We combine MPES with multiple knowledge editing methods including MEMIT Meng et al.
+(2022b), RECT Gu et al.
+(2024b), PRUNE Ma et al.
+(2025) and AlphaEdit Fang et al.
 (2024).
-More details about experimental details can be found in Appendix J.
-4 Overfitting during Knowledge Editing
-As discussed in section 2, locate-then-edit methods can be seen as a two-step fine tuning process.
-The first step uses gradient descent to create intermediate target representations.
-The gradient descent step minimizes average cross-entropy loss for predicting the target fact over the query phrase augmented by ‘$N$’ random prefixes.
-The random prefixes are supposed to represent different contexts in which the edited fact can be recalled, thus aiming to create a more general query representation.
-The average cross-entropy loss over all contexts is optimized as shown:
-$L(\theta)=\frac{1}{N}\sum^{j=N}_{j=1}-log\mathcal{P}_{\theta}[o^{*}|x_{j}+p]$ (2)
-Here $p$ represents the query phrase input to the model, $o^{*}$ represents the target fact to be edited into the model and $x_{j}$ represent random prefixes added to the query phrase.
-(a) MEMIT
-(b) AlphaEdit
-(c) EMMET
-Figure 3: Average downstream performance measured over 6 tasks (sec 2) for MEMIT, AlphaEdit and EMMET.
-We see that MPES is able to delay loss of downstream performance for both Llama3-8B and Llama2-7B with additional gains in efficiencey.
-Table 2: Sequential knowledge editing performance after 10,000 edits for different algorithms. We see that using MPES improves editing metrics across all algorithms and models.
-Method Model Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy
-MEMIT Llama2-7B 81.04 64.67 60.95 67.859 442.59
-Llama3-8B 49.68 49.29 51.31 50.078 373.48
-MEMIT Llama2-7B 88.43 70.83 65.86 73.873 542.1
-+ MPES Llama3-8B 65.78 57.58 50.25 57.176 560.78
-AlphaEdit Llama2-7B 61.1 55.86 53.75 56.74 540.92
-Llama3-8B 72.67 63.44 52.9 61.948 465.81
-AlphaEdit Llama2-7B 84.15 74.94 62.87 72.933 583.4
-+ MPES Llama3-8B 88.43 82.08 56.5 72.832 565.36
-EMMET Llama2-7B 94.98 84.05 55.76 74.331 569.24
-Llama3-8B 85.03 75.06 49.08 65.995 567.34
-EMMET Llama2-7B 96.8 88.4 58.41 77.393 584.28
-+ MPES Llama3-8B 92.96 86.12 55.74 74.424 564.82
-We observe that when the gradient descent process stops, the intermediate target representations are over-optimized on the edited facts and make the probability of the target fact unusually high.
-This can be seen in Table 1, where the edited facts are predicted with a probability much closer to 1 compared to the probability with which an unedited model recalls a fact.
-This phenomenon was also observed by Zhang et al.
+We perform 10,000 sequential edits in batches of 100 as described in section 3.
+The editing and downstream evaluation metrics for sequential editing for both regularization methods can be seen in Table 2 and Figure 3 respectively.
+We see that MPES provides substantial gain across all key metrics.
+Table 2 shows that all editing metrics are enhance significantly for all knowledge editing methods when combined with MPES.
+Infact, Table 2 shows that LTI harms the editing metrics for sequential editing when compared to even the respective baselines.
+The downstream performance metrics can be seen in Figure 3.
+We see that knowledge editing methods augmented with MPES (represented by solid lines with diamond dots) are able to maintain their downstream performance for much longer when compared to the corresponding baselines and LTI.
+However, despite these improvements, downstream degradation still occurs after a few thousand edits especially in larger models like Llama3-8B.
+This suggests that MPES alone may not fully address all failure modes, which we explore further in the next section.
+To summarize, with MPES we present a principled way of stopping the gradient-descent step during knowledge editing which results in improved editing performance, delays loss of downstream performance, and makes current knowledge editing methods much faster.
+6 Norm Growth during Sequential Knowledge Editing
+Previous studies show that sequential knowledge editing increases the norm of the edited matrix Gupta et al.
+(2024b); Yang et al.
+(2024); Ma et al.
 (2025).
-More details on the experimental settings can be found in Appendix B.
-As seen in Table 1, the probability with which the unedited GPT2-XL, Llama2-7B, and Llama3-8B models recall a fact are 39%, 52% and 49% respectively.
-However, facts that get edited into the model with algorithms like Alphaedit or EMMET are predicted with average probabilities of 98-99%.
-The vocabulary size of these models is in the tens of thousands and placing such a high probability mass over every edited fact is a result of clear overfitting.
-We hypothesize that this overfitting of edited facts is one of the reasons why edited models lose their ability to perform downstream tasks when continuously edited using knowledge editing methods.
-4.1 MPES and Knowledge Editing
-To overcome this overfitting over a small subset of edited facts, we propose a variant of early stopping called “most-probable early stopping” (MPES).
-Conventionally, early stopping is used during training while monitoring validation loss, where training is halted when the validation loss stops improving.
-In MPES, we stop the gradient descent process in knowledge editing when the target fact becomes the most probable token for all ‘$q$’ query phrases used for optimization.
-While the gradient descent loss continues to go down and the target probability continues to increase, the rank of the predicted word does not change further.
-With MPES, the rank of the predicted word performs the same function as validation loss in conventional early stopping.
-Apart from preventing the model from becoming overly optimized towards a specific target fact, using MPES for halting gradient descent also has two other advantages.
-Firstly, it simplifies monitoring of the gradient descent process and provides a principled approach to stopping gradient descent which is directly tied to the knowledge editing objective, that is, accurately recalling edited facts in a variety of scenarios without overfitting.
-Secondly, by optimally stopping the gradient descent process, we also improve the efficiency of locate-then-edit algorithms.
-MPES leads to a 39-76% reduction in editing time per edit depending on the editing algorithm and model used (Table 1).
-The average downstream performance of AlphaEdit, MEMIT, and EMMET for Llama2-7B and Llama3-8B over 10,000 sequential edits is shown in Figure 3111For task-wise downstream performance, please refer to Appendix H.
-Comparing baseline methods with MPES, we see that MPES shows significant improvements in downstream performance for both AlphaEdit and MEMIT.
-For EMMET, MPES performs on par with the baseline but is much faster, as shown in Table 1.
-This shows that by halting gradient descent at a critical point where the edited fact is recognized but not overly dominant, MPES helps maintain the model’s general abilities for much longer during sequential editing.
-The editing performance using MPES can be seen in Table 2.
-We see that MPES not only maintains the model’s general abilities for longer but also improves editing performance for all algorithms.
-This is because as the model over-optimizes for edited facts, it not only loses its general abilities but also the ability to recall facts edited in previous iterations, resulting in lower editing scores.
-This also verifies our hypothesis that overfitting over edited facts is one of the reasons behind the loss of model’s general ability.
-To summarize, with MPES we present a principled way of stopping the gradient descent step during knowledge editing which results in improved editing performance, delays the loss of downstream performance, and makes current knowledge editing methods much faster.
-(a) Alphaedit (Llama3-8B)
-(b) MEMIT (Llama3-8B)
-Figure 4: Comparison between norm of edited MLP matrices and norm of unedited matrices after 5,000 and 10,000 sequential edits.
-5 Norm Growth during Sequential Knowledge Editing
-Past work has shown that sequential knowledge editing can lead to an increase in the norm of the edited matrix (Gupta et al., 2024b).
-This disproportionate growth of norm can be seen in Figure 1 and 4 when using AlphaEdit and MEMIT on Llama3-8B.
-Both AlphaEdit and MEMIT modify the second MLP matrix of layers 4 to 8.
-To put the norm growth in perspective with the rest of the model, we plot the norms of the edited matrices after 5k and 10k edits along with the norm of other MLP matrices of the model.
-We clearly see the disproportionate growth in the norm for the edited layers in Figure 4, where the norm grows to more than 10 times its original value when using MEMIT, and twice the original norm when using AlphaEdit, while the norm of the matrices before and after the edited layers is not changed.
-The growth of the norm of the edited matrix is a continuous process during knowledge editing.
-As shown in Figure 1, for not even one edit does the norm of the edited matrix remain constant or decrease during editing.
-The norm always increases!
-But is this growing norm harmful to the model? We answer this question by analyzing the residual stream and reveal a very important property about the inner workings of knowledge editing methods.
-5.1 The Secret Mechanics of Knowledge Editing
-Due to residual connections in an LLM, the final output of a model can be written as a summation of the outputs of the individual sub-modules.
-Let $h^{l}$ represent the intermediate hidden state vectors between each decoder layer.
-Then, the computations within a layer of a general decoder-only LLMs proceed as follows:
-[…]
-$a^{l}$ $=\texttt{Att}(f^{l})$ (4)
-[…]
-$m^{l}$ $=W^{l}_{proj}\sigma(W^{l}_{fc}g^{l}+b^{l}_{fc})+b_{proj}$ (6)
-$h^{l}$ $=h^{l-1}+a^{l}+m^{l}$ (7)
-The intermediate hidden vector between each layer, $h^{l}$, is also sometimes referred to as the residual stream.
-LN1 represents the first LayerNorm (or equivalently RMSNorm for Llama models) that acts just before the attention module and LN2 represents the second LayerNorm just before the MLP module.
-Att represents the self-attention module in an LLM whereas the action of a traditional MLP module is written in terms of the individual weight matrices ($W_{fc},W_{proj}$).
+Figure 1 illustrates this issue, highlighting the extreme norm growth in edited layers of Llama3-8B using MEMIT.
+Here, the norms of the edited layers increase by more than 10 times, while the unedited layer remains unchanged (Figure 1(b)).
+Furthermore, norm growth persists continuously during editing, as shown in Figure 1, for not even one edit does the norm remain constant or decrease.
+While the anomalous norm-growth was observed in prior work, they do not explain how it affects the general ability of the model.
+We answer this question by analyzing the residual stream of the model during large-scale knowledge editing.
+6.1 Explaining Loss of Downstream Performance due to Norm-Growth
+To understand the impact of this norm growth, we analyze how residual connections work in decoder-only LLMs.
+The intermediate hidden vector at layer $l$, represented by $h^{l}$, is also sometimes referred to as the residual stream.
+Each decoder-only layer contains one attention and FFN module that feeds directly into the residual stream through residual connections.
+The exact computations happening within transformer-based decoder-only LLMs can be found in Appendix E.
+Let the output of the attention module at layer $l$ be represented by $a^{l}$, and the output of the FFN module be represented by $m^{l}$.
 As the vectors computed in the attention and MLP modules get added back to the residual stream at each layer, the residual stream represents a summation of an increasing number of vectors as we go deeper into the model.
-A non-recursive formula for the output of the transformer just before unembedding is shown below:
+A non-recursive formula for the output of the transformer just before unembedding and final layernorm is shown below:
 $h^{L}=h^{0}+\sum^{i=L}_{i=1}a^{i}+\sum^{i=L}_{i=1}m^{i}$ (3)
 Here, $L$ represents the total number of layers in a model and $h^{L}$ represents the residual stream after the final layer.
-Note that $h^{L}$ is the activation just before the application of the final layernorm and unembedding matrices.
 Thus, the output vector at the final layer is a summation of the outputs of individual self-attention and MLP sub-modules.
-Now, if the norm of the $W_{proj}$ matrix grows as disproportionately as shown in Figures 1 and 4, the norm of the vectors that are produced from those edited MLP sub-modules will also grow.
-As the norm of a few vectors in the summation in equation 8 grows, these vectors will begin to dominate the sum.
-Proof for this is shown in Appendix C.1, where we show that if the norm of a vector in a summation grows, the overall sum effectively tends towards that vector.
-(a) Average Norm Proportion For Unedited Llama3-8B
-(b) Average Norm Proportion for Llama3-8B using Alphaedit
-(c) Average Norm Proportion for Llama3-8B using MEMIT
-(d) Average Norm Proportion for Llama3-8B using ENCORE
-Figure 5: The figure shows the proportion of contribution of activations vectors produced from each sub-module to the residual stream.
-The edited layers are shown in red.
-We see that the influence of the output of the edited layers grows dramatically after contiuous editing.
+Now, if the norm of the edited MLP matrix grows as disproportionately as shown in Figure 1, the norm of the vectors that are produced from those edited MLP sub-modules will also grow.
+This means that the norms of the vectors $m^{l}$ in the summation corresponding to the edited layers will grow substantially.
+As the norm of a few vectors in the summation grows, these vectors will begin to dominate the sum.
+Proof for this is shown in Appendix D.1, where we show that if the norm of a vector in a summation grows, the overall sum effectively tends towards that vector.
 We also show this effect empirically.
-The growing norm of activation vectors produced by edited layers before and after editing for Llama3-8B can be seen in Figure 5.
-Figure 5(a) shows the percentage norm for activation vectors produced from each sub-module of the unedited Llama3-8B model, in comparison to the sum of their norms [^2] This is averaged over 30,000 generated tokens (Appendix C)..
-After editing using AlphaEdit, which edits layers 4-8 in the Llama model for 10,000 sequential edits (edited layers are shown in red color on x-axis), we can see a drastic increase in the percentage norms for vectors produced by the edited layers (Figure 5(b)).
-For example, the norm of vectors produced by layer 8, which accounted for only 1% of the total norm, after editing using AlphaEdit accounts for 7% of the norm.
-This effect is even more drastic for MEMIT as seen in Figure 5(c), where the activation vectors produced by layer 8 account for almost 40% of the total norm, and the vectors produced by all edited layers account for 85% of the total.
-This gives us a crucial insight into the inner workings of locate-then-edit methods.
-Apart from optimizing for the preservation-memorization objective in equation 1, these methods also implicitly increase the norms of the edited matrices, and as a consequence, the norm of the activation vectors produced from those matrices also increases.
-This makes the representation after the final layer ($h^{L}$) be dominated by the output of edited layers, thus giving the edited layers a larger authority over the final representations.
-This allows edited layers to override the information produced from other parts of the model, consequently leading to the success of these knowledge editing methods.
-This finding is central to explaining the inner workings of locate-then-edit methods, and it also explains their successes and failures.
-This kind of “importance hacking” makes the output of the edited layers more dominant in the final output and is a shortcut taken by these methods to achieve edit success.
+The growing norm of activation vectors produced by edited layers after editing Llama3-8B can be seen in Figure 4.
+After editing using MEMIT, which edits layers 4-8 for 10,000 sequential edits (edited layers are shown in red color on the x-axis), we see a drastic increase in the norm of activation vectors produced by edited layers.
+For example, the activation vectors produced by layer 8 account for almost 40% of the total norm, and vectors produced by all edited layers account for 85% of the total.
+To emphasize how extreme this is, the residual stream for Llama-3-8B is made up of a summation of 65 vectorsLlama-3-8B has 32 layers, each contributing two vectros $a^{l}$ and $m^{l}$ to the residual stream along with the input vector., and 4 out of the 65 vectors coming from the edited layers account for 85% of the total norm.
+The norm-contributions for an unedited model can be seen in Figure 9(a) (appendix), where the four edited layers account for less than 4% of the overall norm.
+Note that this effect cannot be mitigated by LayerNorm or RMSNorm, which just normalize the incoming vectors Brody et al.
+(2023), whereas the norm growth of edited layers changes the content of the final hidden representation by making it consist of mostly the output of the edited layers.
+This gives us a crucial insight into why the growing norm of edited matrices leads to a loss of downstream performance.
+As the norms of the edited matrices increase, and as a consequence, the norm of the activation vectors produced from those matrices also increases, the activation vector at the final layer ($h^{L}$) is dominated by the output of edited layers, thus giving the edited layers a larger authority over the final representations.
+This allows edited layers to override the information produced from other parts of the model, possibly also helping make successful edits.
 However, as the norm of the edited matrix continues to grow when we edit it sequentially, suddenly the final representations begin to largely be composed of the outputs of only the edited layers.
-This makes it impossible for the model to account for the information processed from other sub-modules which might be important to perform other downstream or unrelated tasks.
-We hypothesize that this is another cause of the loss of downstream performance.
-Table 3: Editing performance of ENCORE when compared to baseline MEMIT and modifications using MPES and Norm Constraint.
-Method Model Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy
-MEMIT Llama2-7B 81.04 64.67 60.95 67.859 442.59
-Llama3-8B 49.68 49.29 51.31 50.078 373.48
-MEMIT Llama2-7B 88.43 70.83 65.86 73.873 542.1
-+ MPES Llama3-8B 65.78 57.58 50.25 57.17 560.78
-MEMIT Llama2-7B 90.94 81.31 59.73 74.931 539.58
-+ Norm-Const Llama3-8B 85.72 77.08 58.45 71.86 367.46
-ENCORE Llama2-7B 92.57 82.64 60.43 76.043 560.16
-Llama3-8B 88.77 78.19 60.07 73.707 523.61
-5.2 Introducing ENCORE
-In the above discussion, we show how growing norm of edited matrices is detrimental to the functioning of edited models and hypothesize that it is the cause of loss of downstream performance.
-To test our hypothesis, we propose to add an additional term to the preservation-memorization objective to control this norm growth.
-Thus, we augment the MEMIT objective with a norm-constraint:
+This makes it impossible for the model to account for the information processed from other sub-modules, which might be important to perform other downstream or unrelated tasks.
+6.2 Introducing Norm Constraint
+In the above discussion, we show how growing norm of edited matrices is detrimental to the functioning of edited models and hypothesize how it causes loss of downstream performance.
+To test these conclusions, we propose to add an additional term to the editing objective to control this norm growth.
+Thus, we augment the MEMIT objective with a Frobenius norm-constraint:
 $L(\hat{W})=$ $\underbrace{\lambda_{p}\sum^{P}_{i=1}\left\|\hat{W}k^{i}_{0}-W_{0}k^{i}_{0}\right\|^{2}_{2}}_{\text{preservation}}+\underbrace{\sum^{B}_{j=1}\left\|\hat{W}k^{j}_{e}-v^{j}_{e}\right\|^{2}_{2}}_{\text{memorization}}$ (4)
 $\hskip 50.0pt+\underbrace{\lambda_{n}\left\|\hat{W}-W_{0}\right\|^{2}_{F}}_{\text{norm-constraint}}$
-The original weight matrix is represented by $W_{0}$ and $\hat{W}$ represents the edited matrix. The above objective has a closed-form solution as shown below (proof, Appendix D):
+Figure 4: The figure shows the proportion of contribution of activation vectors from each sub-module to the residual stream. Edited layers are shown in red.
+The original weight matrix is represented by $W_{0}$ and $\hat{W}$ represents the edited matrix. The above objective has a closed-form solution as shown below (proof, Appendix F):
 $\hat{W}$ $=W_{0}+\Delta\hskip 10.0pt\text{where}$ (5)
 $\Delta=(V_{1}-W_{0}K_{1})$ $K_{1}^{T}(\lambda_{p}K_{0}K_{0}^{T}+K_{1}K_{1}^{T}+\lambda_{n}I)^{-1}$
-Editing performance for 10,000 sequential edits using norm-constraint can be seen in Table 3.
-Since we add the norm-constraint to the MEMIT objective  [^3] While adding norm-constraint works well with the MEMIT objective, using norm constraint with EMMET and AlphaEdit does not lead to such extensive improvements.
-We discuss this further in Appendix E., we compare the performance of MEMIT and MEMIT+MPES.
-We see a significant improvement in all editing metrics for both Llama2-7B and Llama3-8B.
-The downstream performance can be seen in Figure 6, where norm-constraint sustains downstream performance for much longer compared to MEMIT and MPES.
+Table 3 and Figure 5 show the editing and downstream performance after 10,000 sequential edits using the norm-constraint.
+Compared to baseline methods such as MEMIT, AlphaEdit, as well as regularization approaches such as PRUNE and RECT, Frobenius norm-constraint version of MEMIT (represented by MEMIT + NC) produces consistent and substantial improvements in editing and downstream performance metrics.
+Specifically, we see in Table 3 that MEMIT + NC outperforms all previous models on all editing metrics except for the generation entropy metric.
+This shows that explicitly controlling norm growth is more effective than prior regularization methods like PRUNE and RECT, particularly when editing at scale.
+When looking at downstream performance in Figure 5, MEMIT + NC also outperforms all prior editing methods when looking for Llama-2-7B, and all methods except AlphaEdit for Llama3-8B.
+Moreover, the observed gains in downstream performance support our hypothesis that uncontrolled norm growth in edited layers leads to downstream degradation.
+Method Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy Llama2-7B Llama3-8B Llama2-7B Llama3-8B Llama2-7B Llama3-8B Llama2-7B Llama3-8B Llama2-7B Llama3-8B MEMIT 81.04 49.68 64.67 49.29 60.95 51.31 67.86 50.08 442.59 373.48 PRUNE 70.80 49.38 62.11 49.63 51.86 51.09 60.60 50.02 280.83 340.22 RECT 82.42 63.17 66.84 56.92 67.39 52.89 71.54 57.36 549.35 588.39 AlphaEdit 61.10 72.67 55.86 63.44 53.75 52.90 56.74 61.95 540.92 465.81 AlphaEdit w/o NC 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 434.17 560.92 MEMIT + LTI + NC 54.41 50.76 51.87 48.89 52.05 51.17 52.75 50.25 516.24 575.59 MEMIT + NC 87.89 85.72 79.10 77.08 59.72 58.45 73.59 71.86 517.86 367.46 MEMIT + MPES + NC 91.09 87.54 78.02 77.97 59.78 59.30 74.03 72.97 545.72 536.43
+Table 3: Editing performance of our approach when compared to baseline MEMIT, AlphaEdit and MEMIT regularization method such as PRUNE and RECT.
 (a) Llama2-7B
 (b) Llama3-8B
-Figure 6: Average downstream performance for during sequential editing with ENCORE compared to baseline of MEMIT and addition of MPES and Norm-Constraint (NC).
-(a) Comparison with unedited layers.
-(b) Norm growth as function of edits.
-Figure 7: Growth in norm of edited matrices for ENCORE for Llama3-8B.
-To achieve our goal of performing 10,000 sequential edits without loss of downstream performance, we combine MPES with the norm-constraint objective in equation 9 and present ENCORE - Early stopping and Norm-Constrained Robust knowledge Editing.
-With ENCORE, we bring back the classical knowledge editing objective of MEMIT in a stronger, faster, and more robust way.
-The editing performance for ENCORE can be seen in Table 3 and Figure 6.
-For Llama2-7B, we see that maximum improvement in downstream performance (Figure 6(a)) comes from the norm constraint and ENCORE performs equally well.
-But ENCORE is not only much faster, but also has a better overall editing performance (Table 3).
-For Llama3-8B (Figure 6(b)), we see that the combination of MPES and norm-constraint is crucial for sustained downstream performance.
-ENCORE also controls the growing norm of edited matrices, as shown in Figure 7.
-As a consequence, the activation norms for ENCORE have a very similar distribution to the original model, which is shown in Figure 5(d).
-Thus, ENCORE is able to perform large-scale knowledge editing without taking the shortcut of importance hacking, which also prevents the loss of downstream performance.
-With ENCORE, we combine the insights presented in this paper about overfitting and implicit importance hacking in locate-then-edit methods and enable long-term sequential knowledge editing.
-ENCORE is able to consistently sustain downstream evaluation metrics of the original model while going through 10,000 sequential edits.
-ENCORE is also 61% faster than MEMIT and 64% faster than AlphaEdit for Llama3-8B [^4] ENCORE takes 1.01 seconds per edit for GPT2-XL, 3.33 seconds for Llama2-7B and 3.35 seconds for Llama3-8B..
+Figure 5: Average downstream performance for during sequential editing with compared to baseline of MEMIT and addition of MPES and Norm-Constraint (NC).
+Additionally, we want to note that AlphaEdit inherently has an explicit norm-constraint in its optimization objective.
+To present the importance of norm-constraint during large scale editing, we remove this term from the objective and evaluate the method, as shown in Table 3.
+To our surprise, we find that AlphaEdit completely fails without the norm constraint, as shown in the row ‘AlphaEdit w/o NC’ in Table 3 .
+Specifically, after 10,000 edits, the model effectively collapes, where Llama2-7B outputs <unk> and Llama3-8B loops on "!" and all editing scores drop to zero.
+We also illustrate the unbound growth of norm in Figures 26(a) and 26(b) in Appendix L.
+These results support our analysis in section 6.1 and re-emphasize that managing norm-growth is essential for stable and successful large-scale editing.
+6.3 Combining MPES and Norm Constraint
+In this section, we combine MPES, which intervenes in the gradient-descent step, with explicit Frobenius norm constraint, which regularizes the weight-update step in MEMIT.
+This combined approach, MEMIT + MPES + NC, achieves the strongest results across both editing score and downstream metrics.
+As shown in Table 3, it consistently improves the consistency of editing scores between models, with the largest gains observed on Llama3-8B.
+More importantly, Figures 5(a) and 5(b) show that the combination of the two regularization methods beautifully preserves downstream task performance with great effectiveness when compared to other algorithms, even after 10,000 sequential edits.
+These findings suggest that MPES and norm-constraint address complementary failure modes during large-scale sequential editing.
+Their combination provides regularization methods that enable scalable knowledge editing while having minimal loss of downstream task performance.
+The generalizability of these findings can be seen by the compatibility of MPES with other knowledge editing methods, and the neccesity and importance of managing norm-growth in both MEMIT and AlphaEdit.
 7 Conclusion
-In this paper we present two major drawbacks of existing knowledge editing methods - (i) overfitting on edited facts and (ii) continuous and disproportionate norm-growth of edited matrix.
-We also present insights into the inner workings of locate-then-edit methods and show that they achieve edit success using the short-cut of norm-increase, which increases the importance of the output of edited matrices.
-We then present ENCORE, a faster and stronger knowledge editing method compared to its predecessors.
-ENCORE is able to perform 10,000 sequential knowledge edits without noticeable loss of downstream performance in about 40% of the time taken by other algorithms for Llama3-8B.
-While many recent works have shown limitations of knowledge editing methods at scale, we show that with a better understanding and appropriate regularizations, these methods can indeed be scaled.
-We hope this brings further research and excitement into the field of knowledge editing.
-Impact Statement
-This paper advances the field of knowledge editing in LLMs by enabling scalable, updates to a model’s stored knowledge without significant model degradation.
-Such techniques could be beneficial for multiple practical and societal uses, including quick updating of time-sensitive facts such as changing a model’s knowledge following natural disasters or newly published research.
-However, as with any technology, there are potential risks with this method.
-The ability to efficiently overwrite and retain specific facts raises concerns about malicious editing or the injection of wrong information.
-Further research on robust governance and the use of these frameworks is still needed.
-Overall, we believe this research can lead to more responsive, up-to-date, and ethically aligned LLMs provided that careful attention is paid to responsible use.
+In this paper, we show that existing knowledge editing methods require appropriate regularization when scaled sequentially to a large number of edits.
+Specifically, we diagnose and address two core failure modes in large-scale knowledge editing, (1) over-optimization of intermediate activations and (2) uncontrolled norm growth of edited matrices.
+To mitigate these, we introduce stage-specific regularization techniques in the form of “most-probable earlys stopping” (MPES) and Frobenius Norm-Constraint, which generalize to multiple knowledge editing methods.
+With the combination of both these regularization methods, we are able to make 10,000 sequential edits while consistently maintaining the downstream performance levels of the unedited model.
+8 Limitation
+Although our method significantly improves editing speed, scaling beyond 10,000 edits still remains challenging due to compute constraints.
+Performing 10,000 sequential edits on Llama3-8B already required a full day of GPU time on our hardware.
+This is because running these edits is combined with a lot of analysis, including downstream performance evaluation and measuring editing accuracy metrics on past edited facts.
+In contrast, baseline methods such as MEMIT combined with LTI were substantially slower, making larger-scale comparisons infeasible.
+Future work could explore more efficient optimization techniques or hardware acceleration to push beyond this limit.
+Acknowledgements
+This work was supported in part by the NVIDIA Academic Grant Program award.
 Appendix A Knowledge Editing Metrics
@@
 GE is computed as the weighted average of bi-gram and tri-gram entropies in the text generated by the edited model.
-A lower GE indicates repetitive text generation, a common failure mode (Meng et al., 2022a; Gupta et al., 2024b).
+A lower GE indicates repetitive text generation, a common failure mode Meng et al.
+(2022a); Gupta et al.
+(2024b).
   5. 5.
-    Score (S): introduced by (Meng et al., 2022a), this composite metric combines edit success, generalization, and locality into a single score.
+    Score (S): introduced by Meng et al.
+(2022a), this composite metric combines edit success, generalization, and locality into a single score.
 It is calculated as the harmonic mean of the Efficacy Score (ES), Paraphrase Score (PS), and Neighborhood Score (NS).
@@
   3. 3.
+    Edited fact probability with LTI - Here we also evaluate the average probability with which a model recalls an edited fact. In this case we used LTI during editing the fact.
+  4. 4.
     Edited fact probability WITH MPES - Here we also evaluate the average probability by which a model recalls an edited fact. In this case, MPES is used during editing the fact.
 In each of the experiments we performed, we passed average numbers over 1000 edited facts with a batch size of 1.
-We use the CounterFact dataset (Meng et al., 2022a) for all of these experiments.
+We use the CounterFact dataset Meng et al.
+(2022a) for all of these experiments.
 We show the result in table 1.
@@
 MPES brings the probability of fact recall for edited facts down to a more natural value, which prevents the overfitting problem that we present in this paper.
-Appendix C Experimental Details on Activation Norm Growth
+Appendix C GPT2-XL Result
+In this section we show the result for GPT2-XL on CounterFact dataset in Tables 4 and 5 and also the downstream average downstream performance in Figure 6.
+As we can see that the result for GPT2-XL also improves the results as well.
+Method Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy MEMIT 94.04 79.91 57.90 74.22 517.37 MEMIT + LTI 82.03 70.51 56.29 67.97 508.35 MEMIT + MPES 92.17 77.01 60.38 72.26 523.57 PRUNE 61.05 58.05 50.00 55.96 579.69 PRUNE + LTI 80.06 69.77 56.07 67.18 542.41 PRUNE + MPES 76.91 59.11 65.96 66.54 563.57 RECT 51.40 49.83 52.17 51.12 409.42 RECT + LTI 47.51 47.37 52.74 49.08 185.96 RECT + MPES 51.99 49.93 54.29 52.01 523.63 AlphaEdit 88.58 70.33 56.04 69.20 580.27 AlphaEdit + LTI 74.20 59.90 54.86 61.98 596.47 AlphaEdit + MPES 95.52 82.08 60.03 76.32 565.44
+Table 4: Knowledge editing performance for GPT2-XL on the CounterFact dataset for different algorithms in combination with MPES.
+Method Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy MEMIT 94.04 79.91 57.90 74.22 517.37 PRUNE 61.05 58.05 50.00 55.96 579.69 RECT 51.40 49.83 52.17 51.12 409.42 AlphaEdit 88.58 70.33 56.04 69.20 580.27 AlphaEdit + w/o NC 47.37 47.37 52.63 49.00 573.40 MEMIT + LTI + NC 84.15 73.12 54.91 68.54 554.47 MEMIT + NC 93.89 80.9 58.00 74.53 504.68 MEMIT + MPES + NC 93.99 79.79 59.98 75.29 517.44
+Table 5: Knowledge editing performance for GPT2-XL on the CounterFact dataset for different algorithms in combination with MPES and NC.
+(a) GPT2-XL downstream performance comparison with MPES
+(b) GPT2-XL downstream performance comparison with MPES + NC
+Figure 6: Average downstream performance for during sequential editing with compared to baseline of MEMIT and addition of MPES and Norm-Constraint (NC).
+Appendix D Experimental Details on Activation Norm Growth
 To assess the impact of the activation vectors generated by the edited layers before and after editing, we conducted an experiment where we edited our model on a total of 10,000 facts using a batch size of 100.
@@
 For each of the 30,000 examples, we calculated the proportion of the activation norm at each layer.
-We then plotted the mean and standard deviation of these proportions for both the edited and unedited models in figures 8 - 13.
+We then plotted the mean and standard deviation of these proportions for both the edited and unedited models in figures 7 - 12.
 As shown in our results, the proportion of activation norms for the layers that were edited is significantly higher than their neighboring layers.
@@
 The edited layers are highlighted with red color on the x-axis.
-C.1 Theoretical Analaysis of Growth of Vector Norm in a Summation
+D.1 Theoretical Analaysis of Growth of Vector Norm in a Summation
 We want to understand the effect of excessive increase in the norm of a vector in a sum of vectors.
-First, let’s start with an easy example where we have a summation of two vectors, $\mathbb{s}=\mathbb{a}+\mathbb{b}$ and then there is excessive increase in the norm of the first vector, that is $\mathbb{s}=k\mathbb{a}+\mathbb{b}$ where $k$ is some positive scalar.
-To evaluate the effect of this increase, we analyses the tendencies of the sum $\mathbb{s}$ as k increases.
-We first want to understand the norm of $||\mathbb{s}||$. We have the following :
-$||\mathbb{s}||^{2}=||k\mathbb{a}+\mathbb{b}||^{2}=k^{2}||\mathbb{a}||^{2}+2k\mathbb{a}\cdot\mathbb{b}+||\mathbb{b}||^{2}$
+First, let’s start with an easy example where we have a summation of two vectors, $\mathbf{s}=\mathbf{a}+\mathbf{b}$ and then there is excessive increase in the norm of the first vector, that is $\mathbf{s}=k\mathbf{a}+\mathbf{b}$ where $k$ is some positive scalar.
+To evaluate the effect of this increase, we analyses the tendencies of the sum $\mathbf{s}$ as $k$ increases.
+We first want to understand the norm of $||\mathbf{s}||$. We have the following :
+$||\mathbf{s}||^{2}=||k\mathbf{a}+\mathbf{b}||^{2}=k^{2}||\mathbf{a}||^{2}+2k\mathbf{a}\cdot\mathbf{b}+||\mathbf{b}||^{2}$
 From this we can clearly see that as $k$ increases the first term quadratic in $k$ will dominate.
-This means that as $k\to\infty$, $||\mathbb{s}||^{2}\to k^{2}||\mathbb{a}||^{2}$, or $||\mathbb{s}||\to k||\mathbb{a}||$, which is the norm of the new vector.
+This means that as $k\to\infty$, $||\mathbf{s}||^{2}\to k^{2}||\mathbf{a}||^{2}$, or $||\mathbf{s}||\to k||\mathbf{a}||$, which is the norm of the new vector.
 Thus, as the norm of one of the vectors in the summation increases, the norm of the summation tends to the norm of that vector with increasing norm.
 Next, we look at the tendencies of the orientation of the summation as the norm of one vector increases.
-Let $\theta$ be the angle between $\mathbb{s}$ and $k\mathbb{a}$.
-Then, the cosine of the angle between the summation and the new vector $k\mathbb{a}$ is as follows (note that angle between $\mathbb{s}$ and $k\mathbb{a}$ is same as the angle between $\mathbb{s}$ and $\mathbb{a}$):
-$\cos{\theta}=\frac{\mathbb{s}\cdot\mathbb{a}}{||\mathbb{s}||||\mathbb{a}||}=\frac{(k\mathbb{a}+\mathbb{b})\cdot\mathbb{a}}{||k\mathbb{a}+\mathbb{b}||||\mathbb{a}||}$
-In the limit of $k\to\infty$, $||\mathbb{s}||\to k||\mathbb{a}||$ as shown above. Thus, the cosine expression in the limit can be written as:
-$\underset{k\to\infty}{\cos{\theta}}=\frac{(k\mathbb{a}+\mathbb{b})\cdot\mathbb{a}}{k||\mathbb{a}||^{2}}=\frac{k\mathbb{a}\cdot\mathbb{a}}{k||\mathbb{a}||^{2}}+\frac{\mathbb{b}\cdot\mathbb{a}}{k||\mathbb{a}||^{2}}$
-$=1+\frac{\mathbb{b}\cdot\mathbb{a}}{k||\mathbb{a}||^{2}}$
+Let $\theta$ be the angle between $\mathbf{s}$ and $k\mathbf{a}$.
+Then, the cosine of the angle between the summation and the new vector $k\mathbf{a}$ is as follows (note that angle between $\mathbf{s}$ and $k\mathbf{a}$ is same as the angle between $\mathbf{s}$ and $\mathbf{a}$):
+$\cos{\theta}=\frac{\mathbf{s}\cdot\mathbf{a}}{||\mathbf{s}||||\mathbf{a}||}=\frac{(k\mathbf{a}+\mathbf{b})\cdot\mathbf{a}}{||k\mathbf{a}+\mathbf{b}||||\mathbf{a}||}$
+In the limit of $k\to\infty$, $||\mathbf{s}||\to k||\mathbf{a}||$ as shown above. Thus, the cosine expression in the limit can be written as:
+$\underset{k\to\infty}{\cos{\theta}}=\frac{(k\mathbf{a}+\mathbf{b})\cdot\mathbf{a}}{k||\mathbf{a}||^{2}}=\frac{k\mathbf{a}\cdot\mathbf{a}}{k||\mathbf{a}||^{2}}+\frac{\mathbf{b}\cdot\mathbf{a}}{k||\mathbf{a}||^{2}}$
+$=1+\frac{\mathbf{b}\cdot\mathbf{a}}{k||\mathbf{a}||^{2}}$
 Thus, as $k\to\infty$, the cosine of angle between the sum and the vector tends to 1, or the angle between the summation and the vector tends to zero.
 This shows that as the norm of a vector in the summation continues to increase, the both the norm and the orientation of the summation tends towards the vector with increasing norm.
-Finally, the same proof can be generalied to a summation of multiple vectors, where $\mathbb{b}$ represents the sum of the remaining vectors.
+Finally, the same proof can be generalied to a summation of multiple vectors, where $\mathbf{b}$ represents the sum of the remaining vectors.
 (a) Average Norm Proportion For Unedited GPT2-XL
 (b) Average Norm Proportion for GPT2-XL using Alphaedit
-(c) Average Norm Proportion for GPT2-XL using EMMET
-(d) Average Norm Proportion for GPT2-XL using MEMIT
+(c) Average Norm Proportion for GPT2-XL using MEMIT
 Figure 7: Activation norm growth for GPT2-XL.
@@
 (b) Average Norm Proportion for Llama2-7B using Alphaedit
-(c) Average Norm Proportion for Llama2-7B using EMMET
-(d) Average Norm Proportion for Llama2-7B using MEMIT
+(c) Average Norm Proportion for Llama2-7B using MEMIT
 Figure 8: Activation norm growth for Llama2-7B.
@@
 (b) Average Norm Proportion for Llama3-8B using Alphaedit
-(c) Average Norm Proportion for Llama3-8B using EMMET
-(d) Average Norm Proportion for Llama3-8B using MEMIT
+(c) Average Norm Proportion for Llama3-8B using MEMIT
 Figure 9: Activation norm growth for Llama3-8B.
 (a) Average Norm Proportion For GPT2-XL using MEMIT
-(b) Average Norm Proportion for GPT2-XL using Norm Constraint
-(c) Average Norm Proportion for GPT2-XL using ENCORE
-Figure 11: Activation norm growth for GPT2-XL using Norm Constraint and ENCORE.
+(b) Average Norm Proportion for GPT2-XL using NC
+(c) Average Norm Proportion for GPT2-XL using MPES + NC
+Figure 10: Activation norm growth for GPT2-XL using NC and MPES + NC.
 (a) Average Norm Proportion For Llama2-7B using MEMIT
-(b) Average Norm Proportion for Llama2-7B using Norm Constraint
-(c) Average Norm Proportion for Llama2-7B using ENCORE
-Figure 12: Activation norm growth for Llama2-7B using Norm Constraint and ENCORE.
+(b) Average Norm Proportion for Llama2-7B using NC
+(c) Average Norm Proportion for Llama2-7B using MPES + NC
+Figure 11: Activation norm growth for Llama2-7B using NC and MPES + NC.
 (a) Average Norm Proportion For Llama3-8B using MEMIT
-(b) Average Norm Proportion for Llama3-8B using Norm Constraint
-(c) Average Norm Proportion for Llama3-8B using ENCORE
-Figure 13: Activation norm growth for Llama3-8B using Norm Constraint and ENCORE.
-Appendix D Proof for ENCORE Objective
+(b) Average Norm Proportion for Llama3-8B using NC
+(c) Average Norm Proportion for Llama3-8B using MPES + NC
+Figure 12: Activation norm growth for Llama3-8B using NC and MPES + NC.
+Appendix E Internal Computations within Decoder-Only LLMs
+The internal computations happening within a transformer-based decoder-only LLM is shown below:
+[…]
+$a^{l}$ $=\texttt{Att}(f^{l})$ (7)
+[…]
+$m^{l}$ $=W^{l}_{proj}\sigma(W^{l}_{fc}g^{l}+b^{l}_{fc})+b_{proj}$ (9)
+$h^{l}$ $=h^{l-1}+a^{l}+m^{l}$ (10)
+The intermediate hidden vector between each layer, $h^{l}$, is also sometimes referred to as the residual stream.
+LN1 represents the first LayerNorm (or equivalently RMSNorm for Llama models) that acts just before the attention module and LN2 represents the second LayerNorm just before the MLP module.
+Att represents the self-attention module in an LLMs whereas the action of a traditional MLP module is written in terms of the individual weight matrices ($W_{fc},W_{proj}$).
+As the vectors computed in the attention and MLP modules get added back to the residual stream at each layer, the residual stream represents a summation of an increasing number of vectors as we go deeper into the model.
+Appendix F Proof for MEMIT + MPES + NC Objective
 First we have that we can write the equation 4 in term of matrix form where we can stack the $k_{0}^{i}$.
@@
 $\Delta=(V_{1}-W_{0}K_{1})K_{1}^{T}(\lambda_{p}K_{0}K_{0}^{T}+K_{1}K_{1}^{T}+\lambda_{n}I)^{-1}$
-Table 5: Editing performance for EMMET with Llama2-7B with norm constraint.
-Method Model Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy
-EMMET baseline Llama2-7B 93.85 87.32 58.07 76.281 579.79
-EMMET best-editing Llama2-7B 94.24 87.2 54.36 74.12 566.08
-EMMET best-downstream Llama2-7B 90.35 85.97 53.65 72.57 541.16
-Appendix E Norm-Constraint with Other Objectives
-While theoretically norm-constraint can be added to both EMMET and AlphaEdit objectives, it doesn’t work well with these objectives.
-AlphaEdit already has a norm constraint in its objective which is still not able to alleviate norm growth as shown in the main paper.
-Moreover, when norm-constraint is used with the EMMET objective, we see no appreciable improvements in downstream performance but incur a penalty in editing scores.
-We believe that this is the nature of how norm-constraint works.
-When the norm-constraint is added, it penalizes change in weights from the original weight values.
-One way to prevent loss of downstream performance is to set the hyperparameter for norm-constraint very high, in which case we will never lose downstream performance.
-But in this scenario, we will also not be able to edit the model.
-In the other case, if we don’t restrict norm growth, the editor becomes free to increase the norm of edited matrices, which causes anomalous behaviors as discussed in the main paper.
-Thus, the norm constraint always comes at the trade-off of editability.
-For the MEMIT objective, the norm-constraint helps both editing and downstream performance and the algorithm is best able to make use of this constraint.
-However, this is not the case for other algorithms.
-The theoretical reasons why this happens are beyond the scope of this work.
-But we provide evidence for this trade-off empirically for EMMET below.
-Table 5 shows the editing performance of the EMMET baseline without any norm-constraint compared against two cases where norm-constraint is added to its objective.
-In one case, we present the best editing score that can be achieved when a non-zero norm-constrain hyperparameter is added to the EMMET objective ($\lambda_{n}=20$).
-In this case, we clearly see that even the best editing score with norm constraint is worse than the baseline EMMET.
-Figure 14(b) shows the downstream performance for this case, which is even worse than the baseline (Figure 14(a)).
-Secondly, in Table 5 we also show the editing score for the hyperparameter that achieves the best downstream performance with non-zero norm-constrain hyperparameter in the EMMET objective ($\lambda_{n}=100$).
-We clearly see a significant drop from baseline editing performance across all metrics.
-The downstream performance of this case is shown in Figure 14(c), which is not much different from the baseline (Figure 14(a)).
-Thus, we show that the norm-constraint does not work well with the EMMET objective.
-(a) EMMET baseline
-(b) Downstream performance for best editing score with norm constraint.
-(c) Downstream peformance for best the scenario where we get best downstream performance with norm constraint.
-Figure 14: Comparison between the effect of adding norm-constrain the EMMET objective. We see no appreciable improvement in downstream performace but incur loss in editing performance.
-Appendix F Editing Hyperparameters and Hardware Details
-Here we present the hyperparameters that we find and used for the MPES, Norm constraint, and ENCORE.
+Appendix G Editing Hyperparameters and Hardware Details
+Here we present the hyperparameters that we find and used for the MPES, Norm constraint, and MPES + Norm constraint.
 Tables 6, 7 and 8 show the hyperparameters for CounterFact dataset and tables 9,10 and 11 show the hyperparameters for the zsRE dataset.
@@
 All experiments in this paper are done on NVIDIA A6000, including experiments where editing speeds of different methods are timed.
+Method Model $\lambda_{p}$ Probability Cut Off AlphaEdit + MPES GPT2-XL 20,000 +1 Llama2-7B 15,000 +0 Llama3-8B 15,000 +0 MEMIT + MPES GPT2-XL 20,000 +2 Llama2-7B 15,000 +1 Llama3-8B 15,000 +2
 Table 6: Hyperparameters for different algorithms with MPES on CouterFact dataset
-Method Model $\lambda_{p}$ Probability Cut Off
-EMMET + MPES GPT2-XL 10,000 +1
-Llama2-7B 15,000 +0
-Llama3-8B 15,000 +0
-AlphaEdit + MPEs GPT2-XL 20,000 +1
-Llama2-7B 15,000 +0
-Llama3-8B 15,000 +0
-MEMIT + MPEs GPT2-XL 20,000 +2
-Llama2-7B 15,000 +1
-Llama3-8B 15,000 +2
+Method Model $\lambda_{p}$ $\lambda_{n}$ Norm Constraint GPT2-XL 20,000 10 Llama2-7B 15,000 10 Llama3-8B 15,000 20
 Table 7: Hyperparameters for Norm Constraint on CouterFact dataset
-Method Model $\lambda_{p}$ $\lambda_{n}$
-Norm Constraint GPT2-XL 20,000 10
-Llama2-7B 15,000 10
-Llama3-8B 15,000 20
-Table 8: Hyperparameters for ENCORE on CouterFact dataset
-Method Model $\lambda_{p}$ $\lambda_{n}$ Probability Cut Off
-ENCORE GPT2-XL 20,000 10 +3
-Llama2-7B 15,000 10 +2
-Llama3-8B 15,000 20 +1
+Method Model $\lambda_{p}$ $\lambda_{n}$ Probability Cut Off MPES + NC GPT2-XL 20,000 10 +3 Llama2-7B 15,000 10 +2 Llama3-8B 15,000 20 +1
+Table 8: Hyperparameters for MPES + Norm Constraint on CouterFact dataset
+Method Model $\lambda_{p}$ $\lambda_{n}$ Norm Constraint GPT2-XL 20,000 40 Llama2-7B 15,000 10 Llama3-8B 15,000 20
+Table 9: Hyperparameters for Norm Constraint on zsRE dataset
+Method Model $\lambda_{p}$ Probability Cut Off AlphaEdit + MPEs GPT2-XL 20,000 +1 Llama2-7B 15,000 +1 Llama3-8B 15,000 +3 MEMIT + MPEs GPT2-XL 20,000 +5 Llama2-7B 15,000 +1 Llama3-8B 15,000 +4
 Table 10: Hyperparameters for different algorithms with MPES on zsRE dataset
-Method Model $\lambda_{p}$ Probability Cut Off
-EMMET + MPES GPT2-XL 10,000 +1
-Llama2-7B 15,000 +1
-Llama3-8B 15,000 +2
-AlphaEdit + MPEs GPT2-XL 20,000 +1
-Llama2-7B 15,000 +1
-Llama3-8B 15,000 +3
-MEMIT + MPEs GPT2-XL 20,000 +5
-Llama2-7B 15,000 +1
-Llama3-8B 15,000 +4
-Table 10: Hyperparameters for Norm Constraint on zsRE dataset
-Method Model $\lambda_{p}$ $\lambda_{n}$
-Norm Constraint GPT2-XL 20,000 40
-Llama2-7B 15,000 10
-Llama3-8B 15,000 20
-Table 11: Hyperparameters for ENCORE on zsRE dataset
-Method Model $\lambda_{p}$ $\lambda_{n}$ Probability Cut Off
-ENCORE GPT2-XL 20,000 40 +4
-Llama2-7B 15,000 10 +2
-Llama3-8B 15,000 10 +3
-Appendix G Norm growth Result
-In this section we present the observation of norm growth during editing for different methods (including Norm Constraint and ENCORE) with different models.
-Figures 15 and 16(c) show the result for different methods for GPT2-XL.
-Figures 17 and 18(c) show the result for Llama2-7B.
-Lastly figures 19 and 20(c) show the result for Llama3-8B.
+Method Model $\lambda_{p}$ $\lambda_{n}$ Probability Cut Off MPES + NC GPT2-XL 20,000 40 +4 Llama2-7B 15,000 10 +2 Llama3-8B 15,000 10 +3
+Table 11: Hyperparameters for MEMIT+ Norm Constraint on zsRE dataset
+Appendix H Norm growth Result
+In this section we present the observation of norm growth during editing for different methods (including Norm Constraint and MPES + Norm Constraint) with different models.
+Figures 13 and 14 show the result for different methods for GPT2-XL.
+Figures 15 and 16 show the result for Llama2-7B.
+Lastly figures 17 and 18 show the result for Llama3-8B.
 (a) Norm growth of GPT2-XL using Alphaedit
-(b) Norm growth of GPT2-XL using EMMET
-(c) Norm growth of GPT2-XL using MEMIT
-Figure 15: Norm growth of GPT2-XL across different methods
-(a) Norm growth of GPT2-XL using MEMIT
-(b) Norm growth of GPT2-XL using Norm Constraint
-(c) Norm growth of GPT2-XL using ENCORE
-Figure 16: Norm growth of GPT2-XL using Norm Constraint and ENCORE
+(b) Norm growth of GPT2-XL using MEMIT
+Figure 13: Norm growth of GPT2-XL using AlphaEdit and MEMIT
+(a) Norm growth of GPT2-XL using NC
+(b) Norm growth of GPT2-XL using MPES + NC
+Figure 14: Norm growth of GPT2-XL using NC and MPES + NC
 (a) Norm growth of Llama2-7B using Alphaedit
-(b) Norm growth of Llama2-7B using EMMET
-(c) Norm growth of Llama2-7B using MEMIT
-Figure 17: Norm growth of Llama2-7B across different methods
-(a) Norm growth of Llama2-7B using MEMIT
-(b) Norm growth of Llama2-7B using Norm Constraint
-(c) Norm growth of Llama2-7B using ENCORE
-Figure 18: Norm growth of Llama2-7B using Norm Constraint and ENCORE
+(b) Norm growth of Llama2-7B using MEMIT
+Figure 15: Norm growth of Llama2-7B using AlphaEdit and MEMIT
+(a) Norm growth of Llama2-7B using NC
+(b) Norm growth of Llama2-7B using MPES + NC
+Figure 16: Norm growth of Llama2-7B using NC and MPES + NC
 (a) Norm growth of Llama3-8B using Alphaedit
-(b) Norm growth of Llama3-8B using EMMET
-(c) Norm growth of Llama3-8B using MEMIT
+(b) Norm growth of Llama3-8B using MEMIT
 Figure 17: Norm growth of Llama3-8B across different methods
-(a) Norm growth of Llama3-8B using MEMIT
-(b) Norm growth of Llama3-8B using Norm Constraint
-(c) Norm growth of Llama3-8B using ENCORE
-Figure 20: Norm growth of Llama3-8B using Norm Constraint and ENCORE
-Appendix H Downstream Performance
-In this section, we show the downstream performance for each method and also for different algorithms.
-Note that the reason why x -axis shows the number of edits to be from 0 to 100 is because it shows a number of gradient updates (or batch index) instead of a number of facts edited.
-Since editing is done in batches of 100, to achieve 10,000 edits we perform a total of 100 batched updates to the models sequentially.
-Thus, to get the total number of edits finished until a certain point, the x-axis needs to be multiplied by 100, which is the batch size.
-Figures 21(a) - 26(a) show the result for the downstream performance for GPT2-XL on both CounterFact and zsRE datasets.
-Figures 27 - 32 show the result for Llama2-7B on both datasets.
-Lastly, figures 33 - 38 show the result for Llama3-8B on both datasets.
-(a) Downstream Performance for GPT2-XL using AlphaEdit
-(b) Downstream Performance for GPT2-XL using AlphaEdit with MPES
-Figure 21: Downstream Performance for GPT2-XL using AlphaEdit and MPES with CounterFact dataset
-(a) Downstream Performance for GPT2-XL using EMMET
-(b) Downstream Performance for GPT2-XL using EMMET with MPES
-Figure 22: Downstream Performance for GPT2-XL using EMMET and MPES with CounterFact dataset
-(a) Downstream Performance for GPT2-XL using MEMIT
-(b) Downstream Performance for GPT2-XL using MEMIT with MPES
-(c) Downstream Performance for GPT2-XL using MEMIT with Norm Constraint
-(d) Downstream Performance for GPT2-XL using MEMIT with ENCORE
-Figure 23: Downstream Performance for GPT2-XL using MEMIT, MPES, Norm Constraint and ENCORE with CounterFact dataset
-(a) Downstream Performance for GPT2-XL using AlphaEdit
-(b) Downstream Performance for GPT2-XL using AlphaEdit with MPES
-Figure 24: Downstream Performance for GPT2-XL using AlphaEdit and MPES with zsRE dataset
-(a) Downstream Performance for GPT2-XL using EMMET
-(b) Downstream Performance for GPT2-XL using EMMET with MPES
-Figure 25: Downstream Performance for GPT2-XL using EMMET and MPES with zsRE dataset
-(a) Downstream Performance for GPT2-XL using MEMIT
-(b) Downstream Performance for GPT2-XL using MEMIT with MPES
-(c) Downstream Performance for GPT2-XL using MEMIT with Norm Constraint
-(d) Downstream Performance for GPT2-XL using MEMIT with ENCORE
-Figure 26: Downstream Performance for GPT2-XL using MEMIT, MPES, Norm Constraint and ENCORE with zsRE dataset
-(a) Downstream Performance for Llama2-7B using AlphaEdit
-(b) Downstream Performance for Llama2-7B using AlphaEdit with MPES
-Figure 27: Downstream Performance for Llama2-7B using AlphaEdit and MPES with CounterFact dataset
-(a) Downstream Performance for Llama2-7B using EMMET
-(b) Downstream Performance for Llama2-7B using EMMET with MPES
-Figure 28: Downstream Performance for Llama2-7B using EMMET and MPES with CounterFact dataset
-(a) Downstream Performance for Llama2-7B using MEMIT
-(b) Downstream Performance for Llama2-7B using MEMIT with MPES
-(c) Downstream Performance for Llama2-7B using MEMIT with Norm Constraint
-(d) Downstream Performance for Llama2-7B using MEMIT with ENCORE
-Figure 29: Downstream Performance for Llama2-7B using MEMIT, MPES, Norm Constraint and ENCORE with CounterFact dataset
-(a) Downstream Performance for Llama2-7B using AlphaEdit
-(b) Downstream Performance for Llama2-7B using AlphaEdit with MPES
-Figure 30: Downstream Performance for Llama2-7B using AlphaEdit and MPES with zsRE dataset
-(a) Downstream Performance for Llama2-7B using EMMET
-(b) Downstream Performance for Llama2-7B using EMMET with MPES
-Figure 31: Downstream Performance for Llama2-7B using EMMET and MPES with zsRE dataset
-(a) Downstream Performance for Llama2-7B using MEMIT
-(b) Downstream Performance for Llama2-7B using MEMIT with MPES
-(c) Downstream Performance for Llama2-7B using MEMIT with Norm Constraint
-(d) Downstream Performance for Llama2-7B using MEMIT with ENCORE
-Figure 32: Downstream Performance for Llama2-7B using MEMIT, MPES, Norm Constraint and ENCORE with zsRE dataset
-(a) Downstream Performance for Llama3-8B using AlphaEdit
-(b) Downstream Performance for Llama3-8B using AlphaEdit with MPES
-Figure 33: Downstream Performance for Llama3-8B using AlphaEdit and MPES with CounterFact dataset
-(a) Downstream Performance for Llama3-8B using EMMET
-(b) Downstream Performance for Llama3-8B using EMMET with MPES
-Figure 34: Downstream Performance for Llama3-8B using EMMET and MPES with CounterFact dataset
-(a) Downstream Performance for Llama3-8B using MEMIT
-(b) Downstream Performance for Llama3-8B using MEMIT with MPES
-(c) Downstream Performance for Llama3-8B using MEMIT with Norm Constraint
-(d) Downstream Performance for Llama3-8B using MEMIT with ENCORE
-Figure 35: Downstream Performance for Llama3-8B using MEMIT, MPES, Norm Constraint and ENCORE with CounterFact dataset
-(a) Downstream Performance for Llama3-8B using AlphaEdit
-(b) Downstream Performance for Llama3-8B using AlphaEdit with MPES
-Figure 36: Downstream Performance for Llama3-8B using AlphaEdit and MPES with zsRE dataset
-(a) Downstream Performance for Llama3-8B using EMMET
-(b) Downstream Performance for Llama3-8B using EMMET with MPES
-Figure 37: Downstream Performance for Llama3-8B using EMMET and MPES with zsRE dataset
-(a) Downstream Performance for Llama3-8B using MEMIT
-(b) Downstream Performance for Llama3-8B using MEMIT with MPES
-(c) Downstream Performance for Llama3-8B using MEMIT with Norm Constraint
-(d) Downstream Performance for Llama3-8B using MEMIT with ENCORE
-Figure 38: Downstream Performance for Llama3-8B using MEMIT, MPES, Norm Constraint and ENCORE with zsRE dataset
+(a) Norm growth of Llama3-8B using NC
+(b) Norm growth of Llama3-8B using MPES + NC
+Figure 18: Norm growth of Llama3-8B using NC and MPES + NC
 Appendix I Editing Performance on zsRE dataset
 Tables 12 - 14 show the editing scores for the sequential editing experiments on zsRE.
+Model Method Edit Score Paraphrase Score Neighborhood Score GPT2-XL AlphaEdit 42.1 33.61 14.61 AlphaEdit + MPES 54.99 43.18 18.40 MEMIT 74.60 61.77 22.40 MEMIT + MPES 75.09 61.58 23.37 MEMIT + NC 74.51 61.90 23.39 MEMIT + MPES + NC 74.46 61.79 23.41
 Table 12: Editing performance for GPT2-XL on zsre dataset
-Model Method Edit Score Paraphrase Score Neighborhood Score
-GPT2-XL EMMET 39.71 27.17 9.78
-EMMET+MPES 53.55 39.42 16.81
-AlphaEdit 42.1 33.61 14.61
-AlphaEdit + MPES 54.99 43.18 18.4
-MEMIT 74.6 61.77 22.4
-MEMIT + MPES 75.09 61.58 23.37
-Norm-Constraint 74.51 61.9 23.39
-ENCORE 74.46 61.79 23.41
+Model Method Edit Score Paraphrase Score Neighborhood Score Llama2-7B AlphaEdit 83.77 77.12 41.96 AlphaEdit + MPES 83.80 77.64 41.97 MEMIT 79.49 74.29 41.80 MEMIT + MPES 83.01 77.45 44.64 MEMIT + NC 88.73 84.05 47.98 MEMIT + MPES + NC 89.10 84.28 48.51
 Table 13: Editing performance for Llama2-7B on zsre dataset
-Model Method Edit Score Paraphrase Score Neighborhood Score
-Llama2-7B EMMET 75.42 69.69 33.89
-EMMET+MPES 84.07 76.9 41.88
-AlphaEdit 83.77 77.12 41.96
-AlphaEdit + MPES 83.8 77.64 41.97
-MEMIT 79.49 74.29 41.8
-MEMIT + MPES 83.01 77.45 44.64
-Norm-Constraint 88.73 84.05 47.98
-ENCORE 89.1 84.28 48.51
+Model Method Edit Score Paraphrase Score Neighborhood Score Llama3-8B AlphaEdit 89.27 82.19 45.23 AlphaEdit + MPES 93.54 85.93 47.32 MEMIT 96.45 90.30 48.91 MEMIT + MPES 96.85 90.76 47.34 MEMIT + NC 90.40 84.58 49.09 MEMIT + MPES + NC 93.15 86.19 49.81
 Table 14: Editing performance for Llama3-8B on zsre dataset
-Model Method Edit Score Paraphrase Score Neighborhood Score
-Llama3-8B EMMET 96.97 90.96 45.26
-EMMET+MPES 96.75 91.31 46.44
-AlphaEdit 89.27 82.19 45.23
-AlphaEdit + MPES 93.54 85.93 47.32
-MEMIT 96.45 90.3 48.91
-MEMIT + MPES 96.85 90.76 47.34
-Norm-Constraint 90.4 84.58 49.09
-ENCORE 93.15 86.19 49.81
-Appendix J Evaluation of Downstream Performance
+Appendix J Downstream Performance
+Figures 19(a) - 20(b) show the result for the downstream performance for GPT2-XL on both CounterFact and zsRE datasets.
+Figures 21(a) - 22(b) show the result for Llama2-7B on both datasets.
+Lastly, figures 23(a) - 24(b) show the result for Llama3-8B on both datasets.
+(a) AlphaEdit and MPES
+(b) MEMIT, MPES, NC, MPES+NC
+Figure 19: Downstream Performance for GPT2-XL using different editing methods with CounterFact dataset
+(a) AlphaEdit and MPES
+(b) MEMIT, MPES, NC, MPES+NC
+Figure 20: Downstream Performance for GPT2-XL using different editing methods with zsRE dataset
+(a) AlphaEdit and MPES
+(b) MEMIT, MPES, NC, MPES+NC
+Figure 21: Downstream Performance for Llama2-7B using different editing methods with CounterFact dataset
+(a) AlphaEdit and MPES
+(b) MEMIT, MPES, NC, MPES+NC
+Figure 22: Downstream Performance for Llama2-7B using different editing methods with zsRE dataset
+(a) AlphaEdit and MPES
+(b) MEMIT, MPES, NC, MPES+NC
+Figure 23: Downstream Performance for Llama3-8B using different editing methods with CounterFact dataset
+(a) AlphaEdit and MPES
+(b) MEMIT, MPES, NC, MPES+NC
+Figure 24: Downstream Performance for Llama3-8B using different editing methods with zsRE dataset
+Appendix K Norm Decrease From MPES
+In this section we want to highlight one observation that we have is that while MPES does reduce the norm of the edited layer shown in Figure 25 by quite a lot compared to just MEMIT alone in Figure 1(a).
+But we still note that the norm growth is still 3 times higher which is the reason why we think that explicit norm control is still needed.
+Figure 25: The figure shows the Norm-growth as function of number edits in MEMIT Llama3-8B with MPES.
+Appendix L Unbound growth of norm in AlphaEdit
+Here we provide the result if we remove the norm constraint from the objective function of AlphaEdit.
+As we can see, once we remove the norm constraint, the norm growth of AlphaEdit just becomes unbound.
+Note that the reason why the x-axis only goes up to 300 edits is that after that the norm value becomes overflow in Python and it therefore it cannot be plot.
+(a) AlphaEdit with Llama2-7B
+(b) AlphaEdit with Llama3-8B
+Figure 26: The figures show the unbound Norm-growth as function of number edits in AlphaEdit
+Appendix M Batch Size Effect
+For this section we show the result for different batch size we see that our method MPES + NC is still effective for all batch size and for all model
+Method Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B MEMIT 91.97 81.73 50.32 77.67 65.58 49.56 58.00 66.22 50.59 73.19 70.45 50.15 503.98 522.72 281.88 PRUNE 92.76 51.80 48.50 78.43 50.15 49.21 57.25 50.17 51.15 73.17 50.70 49.60 505.02 247.88 297.80 RECT 51.01 51.85 48.94 48.92 50.47 49.98 52.57 49.34 50.14 50.79 50.53 49.68 539.58 323.46 192.70 AlphaEdit 91.24 56.92 57.80 73.71 51.38 56.55 56.39 53.73 49.70 70.99 53.92 54.44 586.77 494.64 430.37 MEMIT + MPES + NC 92.57 90.79 89.71 78.19 79.76 80.48 60.36 59.70 57.63 74.70 74.44 73.31 510.81 555.18 539.91
+Table 15: Editing performance of our approach when compared to baseline MEMIT, AlphaEdit and MEMIT regularization method such as PRUNE and RECT using batchsize 10.
+(a) Downstream Performance for GPT2-XL with batch size 10
+(b) Downstream Performance for Llama2-7B with batch size 10
+(c) Downstream Performance for Llama3-8B with batch size 10
+Figure 27: Different Downstream Performance for different models with batch size 10
+Method Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B MEMIT 94.04 81.04 49.68 79.91 64.67 49.29 57.90 60.95 51.31 74.22 67.86 50.08 517.37 442.59 373.48 PRUNE 61.05 70.80 49.38 58.05 62.11 49.63 50.00 51.86 51.09 55.96 60.60 50.02 579.69 280.83 340.22 RECT 51.40 82.42 63.17 49.83 66.84 56.92 52.17 67.39 52.89 51.12 71.54 57.36 409.42 549.35 588.39 AlphaEdit 88.58 61.10 72.67 70.33 55.86 63.44 56.04 53.75 52.90 69.20 56.74 61.95 580.27 540.92 465.81 MEMIT + MPES + NC 93.35 92.57 88.77 78.66 82.64 78.19 59.84 60.43 60.07 74.75 76.04 73.71 523.49 560.16 523.61
+Table 16: Editing performance of our approach when compared to baseline MEMIT, AlphaEdit and MEMIT regularization method such as PRUNE and RECT using batchsize 100.
+(a) Downstream Performance for GPT2-XL with batch size 100
+(b) Downstream Performance for Llama2-7B with batch size 100
+(c) Downstream Performance for Llama3-8B with batch size 100
+Figure 28: Different Downstream Performance for different models with batch size 100
+Method Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B GPT2-XL Llama2-7B Llama3-8B MEMIT 93.55 93.94 74.45 81.31 76.83 61.25 59.66 67.64 57.33 75.47 78.03 63.56 558.93 577.69 457.97 PRUNE 92.42 88.58 62.83 81.27 70.43 54.11 58.79 64.64 52.29 74.75 73.25 56.05 548.27 527.05 461.71 RECT 87.49 87.35 57.36 71.94 60.86 52.31 64.9 73.67 64.34 73.64 72.37 57.59 603.91 566.67 215.44 AlphaEdit 92.61 96.11 91.29 76.78 87.27 76.88 56.09 61.63 68.95 72.03 78.76 78.00 587.19 588.09 593.35 MEMIT + MPES + NC 92.57 93.75 84.29 80.01 80.45 70.51 61.26 65.71 69.48 75.71 78.30 74.19 566.94 593.12 598.86
+Table 17: Editing performance of our approach when compared to baseline MEMIT, AlphaEdit and MEMIT regularization method such as PRUNE and RECT using batchsize 1000.
+(a) Downstream Performance for GPT2-XL with batch size 1000
+(b) Downstream Performance for Llama2-7B with batch size 1000
+(c) Downstream Performance for Llama3-8B with batch size 1000
+Figure 29: Different Downstream Performance for different models with batch size 1000
+Appendix N Evaluation of Downstream Performance
 In this paper, we assess model degradation by measuring downstream performance at regular intervals of edits.
-Our evaluation suite is wide-ranging and consists of the following 6 tasks – sentiment analysis (SST2) (Socher et al., 2013), paraphrase detection (MRPC) (Dolan & Brockett, 2005), natural language inference (NLI, RTE) (Dagan et al., 2005; Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009), linguistic acceptability classification (CoLA) (Warstadt et al., 2019), and massive multitask language understanding (MMLU) (Hendrycks et al., 2020).
+Our evaluation suite is wide-ranging and consists of the following 6 tasks – sentiment analysis (SST2) Socher et al.
+(2013), paraphrase detection (MRPC) Dolan and Brockett (2005), natural language inference (NLI, RTE) Dagan et al.
+(2005); Haim et al.
+(2006); Giampiccolo et al.
+(2007); Bentivogli et al.
+(2009), linguistic acceptability classification (CoLA) Warstadt et al.
+(2019), and massive multitask language understanding (MMLU) Hendrycks et al.
+(2020).
 For each task, we created a subset of 100 examples balanced across all multiple-choice options.
@@
 In order to improve the models’ initial performance and achieve meaningful signals, we provided few-shot examples.
-The few-shot prompt templates used for each task are shown in Figures 39-44.
+The few-shot prompt templates used for each task are shown in Figures 30-35.
 Review : an exhilarating futuristic thriller-noir , minority report twists the best of technology around a gripping story , delivering a riveting , pulse intensifying escapist adventure of the first order